A Selected Bibliography of Publications by, and about, Ettore Majorana

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.42

Title word cross-reference

(3 + 1) [Heß94]. $16.95 [Wea87]. $85.00 [Mag16]. _{3} [BYK^{+}17]. α [Maj06]. e [RI97]. M [Cha00]. Z_{2} [Tol10]. P′ [Bia06, Maj31d, MA06, RI97]. SL(2, C) [BK90].

-bigradings [Tol10]. -branes [Cha00]. -dimensional [Heß94].

0511222 [Esp05a].

1.5 [RI97]. 15.99/$27.50 [Esp10b]. 175pp [Wea87]. 1933 [CCJ^{+}34]. 1938 [Mau09, PR07]. 1976 [Rec78b].

2001 [BD02]. 2006 [Nas07]. 2009 [Seg10]. 20th [APB96, ABP98]. 21st [RCC^{+}06a]. 22 [CCJ^{+}34].

3 [Fio08, Gue21]. 304pp [Esp10b].

4 [Bon09, Zic66].

7 [Tur08].

88 [Fio08]. 88-7438-031-3 [Fio08].

9 [Mag16]. 978 [Bon09, Gue21, Mag16, Tur08]. 978-1-107-04402-9 [Mag16]. 978-88-339-1746-7 [Tur08]. 978-88-7642-331-4 [Bon09]. 978-981-12-0701-3 [Gue21]. 978-981-12-0769-3 [Car20].

a.c [RLF12]. Absorption [DMA07]. Academic [Bon09, RS09, GR08a]. According [MRB75, CGRS07, Mig74, RE98]. achievements [Esp06a]. action [Tol10]. activity [GR08a, Rek83, Bon09, RS09]. ad [Gia83]. Advanced [Zic66, Car20]. Advances [BD02, BD02]. Affair [Eis88, Sci87, Wea87]. after [GHH02, Mai06, Rec98, Rec99a, Rec99c]. Again [Esp05a, Mau09]. age [Mag09, Mag11, SH16, And10, Seg10, Esp10b, Lan17, Nav11, Sla10, Clo10, Dom09]. air [Mau09]. Al [DRM^{+}12]. Alan [RS09]. Albert [Ano05, Kle05]. alcune [Fer27]. algebraic [Var02]. Amaldi [RRR07]. Ammonia [Di 06]. Analysis [DE06c]. analytic [Esp10a]. angular [MC06]. annihilating [Rei12]. annihilation [Dir30]. anomali [Maj31a]. anomalous [Maj31a, MIM06]. Ansatz [Sim06]. Anwendung [Fer28]. Anyons [Kit06]. app [Gue21]. Apparent [And32]. application [Fer28]. approach [EFGP99, Esp06e]. approximant [EFGP99]. après [LL04]. April [BD02]. arbitrario [Maj32c]. arbitrary [Maj32c, MC06]. arbitrio [Fid02]. Archimedes [RS09]. array [SS12]. Article [Man06b, Man06a, Man06c, Rec07b]. Artificial [GR09]. asintotica [Som32]. Aspects [Bon09, RS09, GR08a]. associated [BK90]. Astronomia [RCC^{+}06a]. astronomy [Mig06]. Asymptotic [Som32]. Atomi [Maj32a, Maj32b, Maj29a, Maj31b]. Atomic [Boh21, Fer27, Lan17, PA06, VOB09, BS93, CCJ^{+}34, Fer28, Ing06, LaL93, Mau09, SH16, Tho27]. atomica [Mau09]. atomiques [CCJ^{+}34]. Atoms [BR45, DE04a, Fer28, Mar75, GR06, Maj29a, Maj31b, Maj32a, Maj32b, MI06, MS06b, Spr91]. atormentado [Mag11]. atti [Nas07]. auspices [CCJ^{+}34]. Autoionization [ACM11, VOB09, ACM10b, ACM10a]. autres [Kle05, Ano05]. avoided [DN05]. avventurosa [Tur08].

Basic [Esp10b, Seg10]. Bassani [Fio08]. be [Che12, Fra12, Mau09, Sul87]. beautiful [And10]. believed [Sul87]. Berlin [Nas07]. Berlino [Nas07]. Besprechungen [HDS^{+}98]. Bestandsaufnahme [GHH02]. Bestimmung [Fer28]. Beta [GR09, AEE08, Fio07, Shc86]. Bethe [Goo99]. Bethge [HDS^{+}98]. Bettina [Lan17]. between [BD02, Dra08, Maj31b, Nas07, Led01]. beyond [Kit06]. bias [DRM^{+}12]. bibl [Gue21]. bibliometric [Gin13].

Bibliométrique [Gin13]. bigradings [Tol10]. biografica [Ama66b]. birefringence [Sch03]. Birth [Bt06, Bt10, Lan17, RCC^{+}06a, ACM10b, ACM10a, SH16]. Bohr [vCK17]. Bollati [Tur08]. bomb [LaL93]. Book [Bon09, Car20, Clo10, Dom09, Eis88, Esp10b, Fio08, Gue21, Mag35, Mag16, Nav11, RS09, Seg10, Sla10, Tur08, Wea87, HDS^{+}98]. Books [Esp10b, Seg10]. Boringhieri [Tur08]. borsa [GR12]. branes [Cha00]. breaking [Gue13]. Breakthrough [And10]. Brillant [And10]. Brilliant [Clo10, Dom09, Esp10b, Nav11, Seg10, Sla10, Mag09]. Broglie [Var02]. Bronstein [Ste02]. Brown [HDS^{+}98]. Brussels [CCJ^{+}34]. Bruxelles [CCJ^{+}34]. Buch [HDS^{+}98]. bulk [Spr91].

Calaprice [HDS^{+}98]. calculation [Tho27]. Cambridge [Mag16]. campo [Maj32a, Maj32b]. can [Fra12]. Carcanet [Wea87]. career [Car20, Gin13]. Carlo [RI97]. Carrière [Gin13]. Case [Car20, Gue21, Far07, Nas07, Rec87, RA91, RA00, RA02, Rec08, Rec11, Rei12, Ron11, Sci80, Rec20]. cases [DE06b, DE07a]. caso [Nas07, Rec87, RA91, RA00, RA02, Rec08, Rec11, Ron11]. Catalog [Rec98, Rec99a]. Catania [RCC^{+}06a]. cattedra [Esp09a]. causa [GM28]. caused [GMA06]. Centenary [Bt06, Bt10]. Centennial [RCC^{+}06a, Esp07a]. Centre [Fer76]. Century [ABP98, And10, RCC^{+}06a, APB96]. CERN [Zic66]. CERN-MPI-NATO [Zic66]. certified [DD13]. cesio [GM28]. cesium [BS93, GM28, GMA06]. chain [SS12]. chair [Esp09a]. challenge [CT12]. Chaos [HDS^{+}98]. Chooz [Las07]. Clarifying [GG11]. class [BK90]. Classical [HDS^{+}98]. clock [BS93]. collaboration [Dra06]. Collected [Dir03]. collection [Rec87, RA91, RA02, RA00, Rec08, Rec11]. collisions [RI97]. Colloquium [ACM10b, EF15]. Commemoration [RCC^{+}06a, Ama66a]. comment [Esp05a]. Comments [Fra66, Fra06]. Company [Lan17, RS09]. compelling [Rei12]. complex [Gue13]. Component [Cas06b, ST68, Cas06a]. computing [Che12, Rei12]. Con [Rec87, RA91, RA02, Maj32c]. Concept [HDS^{+}98]. Concerning [Fra66, Fra06]. Conference [RCC^{+}06a, Nas07]. conseil [CCJ^{+}34]. Conservation [GLR14]. Considérations [Hei34, Hei34]. Consistent [Mar75]. constants [ES10]. Contemporary [Lic06, Puc06]. continuitá [DE06a, DE15]. continuity [DE06a, DE15]. Contribution [DE04b]. contributions [ES10]. convegno [Nas07]. Conversation [Goo99]. corrections [GR06, Maj29a]. Correspondence [Gen88]. correzioni [Maj29a]. Corso [Esp05b, Esp04]. Cosmology [HDS^{+}98]. could [Che12]. Course [DE07b, Esp05b, Esp04]. Covariant [Esp98]. Critique [Ano05]. crossing [DN05]. Crystallography [HDS^{+}98].

dal [De 05]. dalla [Rec99c]. Darkness [And10, Clo10, Dom09, Esp10b, Nav11, Seg10, Sla10, Mag09]. data [Rek83]. day [D’A98, Rus90, Rus97]. Death [GR13, Rec99c]. Decay [GR09, AEE08, Fio07, Shc86, Sul87]. deceased [Ron13]. decision [Ron13]. decisione [Ron13]. défient [CT12]. Deflectable [And32]. degli [De 05, Gia83]. Del [Tur08, Esp04, Esp05b, Fid02, GM28, Maj37, Nas07, Rec75a, Ron11]. dell [Maj37]. Della [RS09, De 05, DE06a, DE15, GR12, Mau09]. dell’allieva [CGRS07]. dell’atomo [Fer27]. dell’elettrone [GM28]. dell’elio [Maj31a]. dell’equazione [Som32]. dello [Maj31c]. departed [Kle00]. desaparecimento [Mag11]. desaparición [Bon05, Sci78]. destino [Mau09]. destiny [Zas06a, Zas06b]. detection [Sch03]. Determination [Fer27, Fer28]. determinazione [Fer27]. d’Ettore [Bon13, Gin13]. deuxiéeme [Bon04]. device [Ali10]. devices [MZF^{+}12]. differential [Som32]. differenziale [Som32]. Diffusometry [HDS^{+}98]. dimensional [Heß94]. Dipartimento [RCC^{+}06a]. Dirac [Ano08, Esp12, Fio06, Gar84, GG11, GR87, MRB75, Mig74, zXZ06]. Dirac- [Gar84]. Dirac-Like [MRB75, Mig74]. Disappearance [And10, Clo10, Dom09, Esp10b, Fra10a, GR13, Nav11, Rec75b, Sci96, SW11, Seg10, Sla10, Bel75, Bon04, Bon05, Bon10, Esp10a, GHH02, Hol09, Mag09, Mag11, Rec75a, Rec98, Rec99a, Sci75, Sci77b, Sci77a, Sci78, Sci82, Sci87, Sci97, Sci01, Sci06, Sci12]. Disappeared [Bon02]. disappears [Mau09]. Discover [Ser12]. Discovery [DE07b, GR09, Rog13, Esp04]. discussions [CCJ^{+}34]. disparition [Bon04, Sci12]. disparu [Kle00]. distinctions [GG11]. do [zXZ06]. documenti [Rec75a, Rec87, RA91, RA00, RA02, Rec08, Rec11]. Documento [Esp05b, Esp04]. Documents [Gue21, Car20, Dra06, Rec75a, Rec87, RA91, RA00, RA02, Rec08, Rec11, Rec20]. Doppelbrechung [Sch03]. dot [SS12]. Double [AEE08, Fio07, Shc86, Las07]. Doubling [WGG12, GM28]. down [WGG12]. dr [GR06]. due [BS93, GM28]. d’un [Kle13].

E. [Dra06, Esp06a, Man06a, Pre06, Rec98, Rec06a]. early [Car20]. Easily [And32]. ed [Rec99c]. edita [Rec99b, Rec99c]. Editorial [Lic06]. Edizioni [Bon09, Fio08, RS09]. Effect [FK08, RLF12]. effects [Gar84]. Eigenschaften [Fer28]. Einführung [HDS^{+}98]. einiger [Fer28]. Einstein [Ano05, HDS^{+}98, Kle05, RRR07, vCK17]. elastic [Gar84]. Electrodynamics [Esp98, Esp07b, Esp07c, Gia85, Esp07d, Var02]. electromagnetic [Dir31]. Electron [EN12, GM28, GMA06, Maj37, Gar84]. electrons [De 05, Dir30, MM06a]. Elementary [Fra66, Fra06]. Elemente [Fer28]. elements [Fer28]. elettrone [Maj37]. elettroni [De 05]. elio [Maj31c]. Elusive [Ano17, Che12]. EMC2006 [RCC^{+}06a]. energia [Gia83]. Energy [BW35, GLR14, Ano08, Che12, Mig06]. English [Man05]. Enrico [Lan17, DE06a, DE06b, DE07a, DS09, DS10, DE15, GR09, Mal10, Seg70, SH16]. Enter [Bro12]. entities [Rei12]. epistolario [Rec87, RA91, RA00, RA02, Rec08, Rec11]. epistomology [Bon13]. Equation [MRB75, Mig74, Sim06, Esp02, Esp12, PW34, Som32]. Equations [Cas06b, Cas06a, vBT85]. era [Mag11]. Erasmo [Car20, Gue21]. Erice [BD02, Rec78b, Zic66]. esclusione [De 05]. Esposito [Mag16]. Essays [APB96, ABP98]. était [Ano05, Kle05]. ethics [Far07]. Ettore [And10, Clo10, Dom09, Esp06g, Esp10b, Fer76, HDS^{+}98, Mag16, Mau09, Nav11, RCC^{+}06a, Rec99a, Sci96, SW11, Seg10, Sla10, Zic66, Zie00, Alu13, Ama66a, Ama66b, AL66, Ama68a, AL06, Ama68b, Ano65, Ano77, Ano13, ACM10b, ACM10a, ACM11, Bt06, Bt10, Bel75, Bon09, Bon13, Cas13, CGRS07, D’A98, DE06a, DE06b, DE07a, DE15, DE04b, DE07b, Dra08, EMvdMR03, Esp04, Esp05b, Esp06b, Esp06c, Esp06d, Esp06h, Esp06g, Esp06f, Esp07a, Esp07b, Esp07c, Esp08, Esp09a, ERvdMB09, Esp09b, Esp10a, ES10, EAW15, Far07, Fid02, Fin02, Gen88, GHH02, Gia83, Gin13, GR05, GR06, GR08a, GR08b, GR10, GR12, GR13, Hol09, Jac06a, Jac06b, Kle00, LaL93, Mag09, Mag11, Mal10, Man05, Man06b, Man06c, Mau09, Meo13, MRB75, Mig74]. Ettore [Nas07, PR07, Puc06, Rec75a, Rec75b, Rec78a, Rec98, RE98, Rec99b, Rec99c, Rec06b, Rec07a, Rec07b, RE07, Rec09, Rek83, RS09, Ron13, Rus90, Rus97, Sci96, SW11, Seg88a, Ste02, Zas06a, Zas06b, Zic06a, Zic06b, Zic06c, Zic07, Zie00, Fio08]. Étude [Gin13]. Euclidean [Bor93, KT90]. Eureka [RS09]. Evidence [Rec75b, Rec78a, Rec87, RA91, RA00, RA02, Rec08, Rec11, Rei12]. evidential [Fra04]. Exact [Sim06]. exactly [Kit06]. examination [Esp10a]. Exchange [BW35, Bri06]. exclusion [De 05]. Existence [And32, Cha32b, Cha00, Cha32a]. Exotic [Heß94]. experiment [Las07]. experimental [Dra06, Dra08]. Experiments [Dra06, DMA07]. expression [GR06, Maj29a]. extraordinária [Mag11]. Extraordinary [And10, Clo10, Dom09, Esp10b, Nav11, Seg10, Sla10, Mag09, Mag11].

F [Fio08]. fact [Fra12]. Fail [GLR14]. Fall [Sci80]. Fano [ACM11]. Faster [zXZ06]. fate [Mau09]. Fault [DD13]. Fault-tolerant [DD13]. Felde [Sch03]. Fermi [Lan17, DE06a, DE06b, DE07a, DS09, DS10, DE15, DE04a, EFGP99, Esp02, Esp05a, GR05, GR08b, GR09, Mal10, Mar75, Seg70, Seg88b, SH16, Som32, Spr91]. Fermion [Ano17, Bro12, Che12]. fermionic [Bor93]. Fermions [FK08, Ali10, Ano12, Bor93, Cha00, DRM^{+}12, DD13, EF15, Fra10b, MZF^{+}12, Rei12, Sta17, WGG12]. fermo [GR12]. ferromagnetism [Esp09b]. Feshbach [Bia06, VOB09]. fiction [Fra12]. field [Dir31, Maj32a, Maj32b, MI06, Sch03]. Fields [BR45, Mar75, ST68, Tho27]. Finding [Esp05b, Sul87]. First [Rei12, Rec78b]. Fisica [DE06a, DE15, Esp04, Esp05b, RCC^{+}06a, Rec07b, Dra08, Fid02, Maj42, ME06, Mau09]. fisico [Rec75a, Tur08]. Fleeting [Esp06b]. flip [Ing06]. Fois [Ano05, Kle05]. Following [DE04b]. force [Bri06]. Forces [Bri65a, HDS^{+}98]. Forgotten [GR08b, GR05]. form [Sul87]. formation [Maj31c, MS06a]. formazione [Maj31c]. Formula [Sch77, DN05]. Formulation [Esp98, Gia85]. försvinnande [Sci77b]. forward [Rei12]. found [Mau09, Sul87]. foundations [Var02]. Four [Esp06c, Esp06d]. fra [Maj31b]. fractional [RLF12]. Francesco [RS09, Bon09]. francese [Esp06g]. Franco [Tur08]. free [Fid02]. French [Alu13, Ano13, Bon04, Bon13, CCJ^{+}34, CT12, Esp06g, Esp07b, Esp07c, Gin13, Hei34, Kle00, Kle05, Kle13, LL04, Sci12]. Frequency [BS93]. Fundamental [ES10].

Gamba [Tur08]. gefundene [Sch03]. geheime [Zie00]. general [Car20, GR06, Maj29a, Hei34]. generale [Maj29a]. générales [Hei34]. generator [DD13, RI97]. génial [Kle00]. genio [Bon02, Mau09, Mag11]. Genius [And10, Bon02, Clo10, Dom09, Esp10b, Nav11, Seg10, Sla10, Zic06b, Zic06c, Esp06b, Esp09a, Kle00, Mag09, Mag11, Mau09, Zic06a, Zic07]. gennaio [Nas07]. Gentile [Gen88]. German [Fer28, GHH02, HDS^{+}98, KK02, Maj33b, PW34, Sch03, Sci80, Sci96, SW11, Zie00]. Gessen [And10]. Gilda [CGRS07]. Gino [Lan17]. giorno [Rus90, Rus97]. Giovanni [Gen88]. Giuseppe [Fio08]. Glimpsed [Ano12]. Grand [Mag11, Mal10]. grande [Mag11]. graphene [DGS09]. Gravitational [DMA07]. Gravity [BD02]. Group [ST68, Wig39, Dir45, Wig89]. Guerra [Bon09, RS09]. Guide [Dra06].

Hans [Goo99]. Hartree [Mar75]. hb [Esp10b]. Heaviside [Sim06]. heavy [RI97]. Heisenberg [Bri65b]. held [BD02, CCJ^{+}34]. helium [Maj31a, Maj31c, MS06a, MIM06]. Henry [Lan17]. heritage [Esp08]. High [Mig06, VOB09]. Hirshfeld [RS09]. historian [Car20]. Historical [ABP98, APB96, Fin02]. history [Fra04]. Hoerlin [Lan17]. Hole [Esp07b, Esp07c]. Holt [Lan17]. Homage [Jac06a, Jac06b]. Hrn [Sch03]. Human [Rec07a, Rec06a]. hybrid [MZF^{+}12]. hydrogen [Maj31b, MS06b].

Idea [Dvo97]. idrogeno [Maj31b]. ihre [Fer28]. II [Dir03]. im [Sch03]. Imaginary [Dvo97]. Impact [Lic06]. improvement [GR05]. improvements [Spr91]. InAs [DRM^{+}12]. Inaugural [PR07, MPR06]. Incomplete [Bia06, Maj31d, MA06]. incompleti [Maj31d]. Induced [GR09]. inedita [Rec99b, Rec99c, Seg88a]. inedite [Gen88]. inedito [Esp06g]. Infinite [Cas06b, ST68, Cas06a]. Infinite-Component [ST68]. Influence [GR09]. INFN [Pet06]. Infrared [Di 06]. ingegno [Esp09a]. Inhomogeneous [Wig39, Wig89]. inquisidor [Mag11]. Inquisitor [Mag11, Mal10]. Inquisitore [Mal10]. institut [CCJ^{+}34]. Institute [BD02, CCJ^{+}34, Zic66]. Insulator [FK08]. integral [BK90, Esp06e]. integrals [Bor93]. integration [Som32]. Integrazione [Som32]. Intellectuals [RRR07]. Intellettuali [RRR07]. intensità [GM28]. intensities [GM28]. intensity [GMA06]. interactions [Zic66]. interdisciplinary [Rec78b]. interest [Car20]. International [Zic66, CCJ^{+}34, Nas07, RCC^{+}06a]. internazionale [Nas07]. Interplay [BD02]. intrinseco [Maj32c]. intrinsic [Maj32c, MC06]. Introduction [HDS^{+}98, Rec06b, Rec09]. Introduzione [Rec09]. Investigation [Di 06]. ion [Maj31c, MS06a]. ione [Maj31c]. ionizzati [Maj29a]. ions [GR06, Maj29a]. irrevocabile [Ron13]. irrevocable [Ron13]. ISBN [Bon09, Car20, Fio08, Gue21, Mag16, Tur08]. Italian [Ama68a, Ama68b, Bel75, Bon10, Bon02, Cas06c, CGRS07, De 05, DE06a, DE15, Dra08, Esp04, Esp05b, Esp06h, Esp06g, Esp09a, Fer27, Fid02, Fin02, GM28, Gen88, GR10, Led01, Maj28, Maj29a, Maj29b, Maj31a, Maj31b, Maj31c, Maj31d, Maj32a, Maj32b, Maj32c, Maj33a, Maj37, Maj42, Maj87, ME06, Mal10, Mau09, Meo13, Nas07, Pon87, Rec75a, Rec87, RA91, Rec99b, Rec99c, RA00, RA02, RRR07, Rec07b, Rec08, Rec09, Rec11, Ron11, Ron13, Rus90, Rus97, Sci75, Sci77a, Sci82, Sci97, Sci01, Sci06, Seg88a, Som32]. Italy [BD02, PR07, RCC^{+}06a].

J [And10]. Jahre [GHH02]. Jan [PR07]. January [Nas07]. Jeune [Kle13]. João [Esp10b, Nav11, Seg10, Sla10]. Jordan [Var02]. Josephson [RLF12]. Jr. [Gen88]. July [Zic66]. June [Zic66].

Kernphysik [HDS^{+}98]. Kerntheorie [Maj33b]. Kimmich [HDS^{+}98]. Knauf [HDS^{+}98]. Kompakt [HDS^{+}98]. Kvantovaya [Dir03].

laboriosa [De 05]. laborious [De 05]. L’affaire [Fra12]. Landau [DN05]. Langkau [HDS^{+}98]. L’articolo [Rec07b]. last [Meo13]. lasting [Zic07]. later [Esp08]. latest [Esp06a]. Law [Sul87]. Laws [Rec07b, Fid02, Maj42, MM06b, Man05, Man06a]. Leben [Zie00]. Lecture [PR07, Esp06h, Esp06f, MPR06]. Lectures [Maj87, ME06]. left [Rec98, Rec99a, RE07]. Legacy [Lan17, RCC^{+}06a, RS09, Esp07a, Puc06, Rec10, Zic07]. legend [Ano13]. légende [Ano13]. Leggi [Rec07b, Fid02, Maj42]. lengths [ES10]. Leonardo [Wea87, Eis88, Far07, Fra10a, Fra12]. L’épistémologie [Bon13]. Letter [Seg88a]. lettera [Seg88a]. letteratura [Nas07]. Lettere [Gen88]. Letters [Gue21, Car20, Dra06, Rec87, RA91, RA00, RA02, Rec08, Rec11, Rec20]. Level [Car20]. lezione [Esp06h]. Lezioni [Maj87, ME06]. liberto [Fid02]. Life [And10, Cas13, Clo10, Dom09, Eis88, Esp10b, Nav11, RS09, Seg10, Sla10, Mag09, Mag11, Rec78a, RE07, Rec09, Rek83, Zie00]. Light [DMA07, Var02]. Like [MRB75, Mig74]. Lindström [HDS^{+}98]. linear [SS12]. lines [GM28, GMA06]. liquid [BYK^{+}17]. Literature [Eis88, Nas07]. Little [Lan17]. livre [Ano05]. Logical [DE06c]. long [Zic07]. long-lasting [Zic07]. L’opera [Ama68a, Rec99c]. Lorentz [Dir45, ST68, Wig39, Wig89]. Low [Ano08, Ing06]. L’ultimo [Meo13].

M [And10]. Magnetic [Ano17, BR45, Maj32a, Maj32b, MI06, Sch03, Sta17]. magnetico [Maj32a, Maj32b]. magnetischen [Sch03]. Magueijo [And10, Esp10b, Nav11, Seg10, Sla10]. mai [Mau09]. MAJOR [RI97]. Majorana [And10, Bon10, Bon09, Car20, Clo10, Dom09, Esp10b, Fer76, Fio08, GHH02, Gue21, KK02, Mag16, Mau09, Nav11, RCC^{+}06a, Rec87, RA91, Rec99a, RA02, Rec08, RS09, Sch03, Sci12, Seg10, Sla10, Zic66, Ali10, Alu13, Ama66a, Ama66b, AL66, Ama68a, AL06, Ama68b, Ano65, Ano77, Ano08, Ano12, Ano13, Ano17, ACM10b, ACM10a, ACM11, AEE08, Bt06, Bt10, BK90, BS93, Bel75, Bia06, Bon04, Bon05, Bon10, Bon02, Bon13, Bor93, BW35, Bri65b, Bri06, Bro12, Cas06b, Cas06a, Cas13, Cas06c, CGRS07, Cha00, Che12, D’A98, DRM^{+}12, DE06a, DE06b, DE07a, DE15, DD13, DN05, DE04a, Di 06, DE08, DGS09, DE04b, DE06c, DE07b, Dra06, DMA07, Dra08, Dun96, Dvo97, Eis88, EF15, Esp98, Esp02, Esp04, Esp05a]. Majorana [Esp05b, Esp06b, Esp06c, Esp06d, Esp06h, Esp06g, Esp06e, Esp06a, Esp06f, Esp07a, Esp07b, Esp07c, Esp07d, Esp08, Esp09a, ERvdMB09, Esp09b, Esp10a, ES10, Esp10b, EN12, Esp12, EAW15, Far07, Fid02, Fin02, Fio06, Fra66, Fra06, Fra10b, Fra13, FK08, Gar84, Gen88, GHH02, Gia83, Gia85, Gin13, GG11, GR87, GR05, GR06, GR08a, GR08b, GR10, GR12, GR13, Heß94, Hol09, HDS^{+}98, Ing06, Jac06a, Jac06b, Kal00, KK02, Kle00, KT90, LaL93, Lic06, Mag09, Mag11, Mai06, Mal10, Man05, Man06a, Man06b, Man06c, Mau09, Meo13, MRB75, Mig74, MZF^{+}12, Nas07, Pav70, Pes06, Pon87, Pre06, PR07, Puc06, PA06, RI97, Rec75a, Rec75b, Rec78a, Rec87, RA91, Rec98, RE98, Rec99b, Rec99c, RA00, RA02, Rec06a, Rec06b, Rec07a]. Majorana [RRR07, Rec07b, RE07, Rec08, Rec09, Rec10, Rec11, Rec20, Rei12, Rek83, RLF12, Ron11, Ron13, Rus90, Rus97, SS12, Sch03, Sch77, Sci75, Sci77b, Sci77a, Sci78, Sci80, Sci82, Sci87, Sci96, Sci97, Sci01, Sci06, SW11, Sci12, Seg88a, Shc86, Sim06, Sta17, Ste02, ST68, Tol10, Var02, VOB09, Wea76, Wea87, Wil09, Wil12, WGG12, zXZ06, Zas06a, Zas06b, Zic06a, Zic06b, Zic06c, Zic07, Zie00, vBT85, Fra10a]. Majorana-neutrino-mass [Gar84]. Majorana-Teilchen [KK02]. Majoranas [Sci77b]. Man [RS09, Ama66a]. manoscritto [Esp06g]. Manuscript [Esp07b, Esp07c, Esp06g]. Manuscripts [RE07, Rec98, Rec99a]. March [D’A98, Mau09, Rus90, Rus97]. marzo [Mau09, Rus90, Rus97]. mass [AEE08, Fio07, Gar84, vBT85]. Match [Lan17]. materialize [Wil12]. Mathematical [And10, Mag16, EAW15]. Matter [Che12, Spr91]. Matvei [Ste02]. May [BD02, Sul87]. meaning [RE98]. Measurement [Fio07]. meccanica [Maj29b]. Mechanics [DE04b, Esp06e, Gue13, Maj29b, Zas06a, Zas06b]. Meeting [CCJ^{+}34]. Memories [Wei85]. Men [Ste02]. Method [Fer27, Fer28]. Methode [Fer28]. metodo [Fer27]. minds [And10]. mine [Sul87]. Missioni [GR12]. misteri [Esp09a]. misterioso [Mag11]. Mistero [Cas06c]. Model [DE04a, GR08b, Esp05a, GR05, Kit06, Tol10]. modes [DGS09, Wil12]. modules [Tol10]. molecolare [Maj31c]. Molecular [PA06, Maj31c, MS06a]. momento [Maj32c]. Momentous [Lan17]. momentum [Maj32c, MC06]. monopoles [Rec78b]. Monte [RI97]. Moreno [Esp04, Esp05b]. Moro [Eis88, Wea87, Sci87, Fra12]. MPI [Zic66]. Mr. [Sch03]. Murder [Wea87]. mysteries [Esp09a]. Mysterious [And10, Clo10, Dom09, Esp10b, Nav11, Seg10, Sla10, Hol09, Mag11]. Mystery [Cas06c, Eis88, Wea87, Zic06b, Zic06c, Esp10b, Zic06a].

nach [GHH02]. Nadia [Bon09, RS09]. Namely [De 05]. nanowire [Ano12, DRM^{+}12, MZF^{+}12, RLF12]. Naples [CGRS07, Maj87, Pre06, PR07]. Napoli [CGRS07, Maj87]. NATO [BD02, Zic66]. Nauchnye [Dir03]. Nauchnykh [Dir03]. navajeros [Sci78]. negativa [Gia83]. nel [Mau09]. nella [Fid02, Maj42, Rec07b]. nelle [Fid02, Maj42, Rec07b]. nell’insegnamento [DE06a, DE15]. neutral [De 05, GR06, Maj29a]. neutri [Maj29a]. Neutrino [Cab06, GLR14, KK02, Las07, Sut92, AEE08, Esp07a, Fio07, Gar84, Mai06, Mig06, Shc86, Var02, Gar84]. neutrino-electron [Gar84]. Neutrinos [Fio06, Pet06, zXZ06, Ano08, AEE08, CT12, Fra04, RI97, vBT85, CT12]. neutro [De 05]. Neutron [BYK^{+}17, Cha32b, GR09, Rog13, Cha32a]. Neutron-Induced [GR09]. Neutrons [Ano17]. never [Mau09]. next [LL04]. NMR [HDS^{+}98]. non [Mau09]. Nonintegrability [HDS^{+}98]. nonorientable [GR87]. Normale [Bon09, RS09]. Nota [Ama66b]. Note [BW35, Lic06]. notebook [Cas13]. Notes [EMvdMR03, RS09, ERvdMB09, MPR06, Spr91]. noyau [Hei34]. noyaux [CCJ^{+}34]. Nuclear [And10, Bri65a, Clo10, DE08, Dom09, Esp10b, HDS^{+}98, Maj65, Nav11, PA06, Seg10, Sla10, EF15, Mag09, Mag11, Maj33b, MR06]. nuclei [CCJ^{+}34, Maj28, Maj29b, Maj33a]. nucleo [De 05]. Nucleus [Maj06, De 05, Hei34]. nulla [Mau09]. number [DD13]. Nuove [Maj33a]. nuovi [Rec75a]. nuovo [Ron11].

o [Mag11, Maj29a, Mag11]. Occasion [Bt06, Bt10]. October [CCJ^{+}34, RCC^{+}06a]. octobre [CCJ^{+}34]. Old [DS09, DS10]. Ombre [Led01]. one [And10]. opera [AL66, AL06, Rec99b]. Operational [Sim06]. opere [Rec09]. Oppenheimer [Gia85, Var02]. optical [GM28, GMA06]. orientati [Maj32a, Maj32b]. Oriented [MI06, Maj32a, Maj32b]. Origin [HDS^{+}98]. Original [Dra06]. oscillations [Cab06, Las07]. Oscillator [Pes06]. ottici [GM28]. outlines [RE07]. Ovvero [De 05, Fid02].

Padé [EFGP99]. pages [RS09]. Panisperna [Tur08, Mal10]. Papa [Mal10]. Paper [Esp05b, Fra66, Fra06, Gue21, Man05, Esp04]. paperback [Car20]. Papers [Bt06, Bt10, Dir03, DE06c, Fio08]. parafermions [Dun96]. particelle [Maj32c]. Particle [Maj06, Ser12, EF15, KK02]. Particles [Fra66, Fra06, Che12, Maj32c, MC06, RLF12]. particolare [Esp06h]. path [Esp06e]. path-integral [Esp06e]. Pauli [Esp07d, Wei85]. peaks [DRM^{+}12]. peculiar [Esp06h, Esp06f]. Pedagogic [Spr91]. People [Wea87]. perception [Fid02]. percezione [Fid02]. Perfect [And10]. perfectly [And10]. periodic [Fer28]. periodischen [Fer28]. Personal [Wei85]. Personality [Rec07a, Rec06a]. Pfaffians [Bor93]. phenomenological [EAW15, Mag16]. phenomenon [Gue13]. Philosophie [Alu13]. philosophy [Alu13]. Photon [MRB75, Mig74, RE98]. Physical [RE98, ES10]. Physicien [Kle13]. physiciens [CT12]. Physicist [Rec75b, Kle13, Rec75a, Seg70]. Physicists [Ser12, CT12]. Physics [ABP98, BD02, DE06b, DS09, DE08, DE06c, DE07b, EMvdMR03, Esp04, Esp05b, Esp06d, Lan17, Lic06, Mag35, Mag16, PR07, PA06, RCC^{+}06a, Rec07b, SH16, Sul87, Zic66, vCK17, APB96, DE06a, DE07a, DS10, DE15, Dra08, EF15, Esp06c, ERvdMB09, EAW15, Fid02, Ing06, Maj42, ME06, MM06b, Man05, Man06a, Mau09, Puc06, Wil12, CCJ^{+}34, Esp05a]. physics/0511222 [Esp05a]. Physik [HDS^{+}98]. physique [CCJ^{+}34]. piece [Ron11]. piroscafo [GR10]. Pisa [Bon09, RS09]. Pope [Lan17, Mal10, SH16]. Portuguese [Mag11]. positivamente [Maj29a]. positive [GR06, Maj29a]. Positives [And32]. positron [Maj37]. positrone [Maj37]. positrons [MM06a]. Possible [Cha32b, MIM06]. postgraduate [Car20]. Potere [RRR07]. Power [RRR07]. pp [Bon09, Car20, Fio08, Gue21, Mag16, Tur08]. pp. [Lan17]. practical [SS12]. Prediction [GLR14]. Preface [RCC^{+}06b]. Preliminary [MPR06]. present [Zic66]. Presentation [Man05]. Press [Mag16]. presumed [Maj31a]. presunti [Maj31a]. Prison [Wea87]. Problem [EN12, DN05, EFGP99]. problema [Gia83]. problems [Zic66]. Proceedings [BD02, Rec78b, Nas07]. promemoria [Ron11]. Properties [Fer27, CCJ^{+}34, Fer28]. proprietá [Fer27]. propriétés [CCJ^{+}34]. proton [De 05]. protone [De 05]. protons [Dir30]. proximate [BYK^{+}17]. Proximity [FK08]. Pseudopolar [MS06b, Maj31b]. pseudopolare [Maj31b]. Publication [GR08b, Gin13, GR05]. Publications [Gin13]. published [Rec99b, Rec99c]. pugnalatori [Sci77a]. puis [LL04]. punto [GR12].

Quanta [vCK17]. Quantised [Dir31]. Quantisierung [PW34]. quantistica [Maj28]. quantization [PW34]. Quantum [BD02, DS09, Dir03, DE04b, Esp06e, Esp07b, Esp07c, Gia85, HDS^{+}98, Wil12, Zas06a, Zas06b, vCK17, BYK^{+}17, Che12, DS10, DD13, Maj28, Rei12, SS12, Var02]. quantum-dot-superconductor [SS12]. Quasi [DE08]. Quasi-Stationary [DE08]. Quirino [Dra08, Dra06, DMA07].

Rabinovitch [Wea87, Eis88]. Race [Fra10b]. Radioactive [Maj06, Maj28, Maj29b]. Radioactivity [GR09, Sul87]. radioattivi [Maj28, Maj29b]. Ragazzo [Tur08]. random [DD13]. random-number [DD13]. Rapporto [RRR07]. Rapports [CCJ^{+}34]. Rare [Sul87]. Rasetti [Tur08]. ray [HDS^{+}98]. Re [Fra12]. Re-writing [Fra12]. reaction [Maj31b, MS06b]. Real [Dvo97, Eis88, Tol10, vCK17]. Realizing [SS12]. really [Fra04]. Reazione [Maj31b]. Recami [Car20, Gue21]. Recollection [Rec98, Rec99a]. Recollections [ABP98, APB96]. refuge [Meo13]. related [Rec78b]. Relationship [RRR07]. relative [Bor93]. Relativistic [MC06, PW34, Maj32c]. relativistica [Maj32c]. relativistischen [PW34]. Relaxometry [HDS^{+}98]. remember [Ama68b, Rec99c]. reminder [Ron11]. rendered [And10]. renewal [DE06a, DE15]. replica [Gue13]. reported [Che12]. Reports [CCJ^{+}34]. representation [BK90]. Representations [Pav70, ST68, Wig39, Dir45, Wig89]. research [Car20, ERvdMB09]. researcher [Esp06a]. researches [Maj33a]. Resonances [VOB09, Bia06]. rest [vBT85]. Résurrection [Kle13, Kle13]. returns [Wil09]. Revelations [Bel75]. Review [Bon09, Car20, Clo10, Dom09, Eis88, Fio08, Gue21, Mag16, Nav11, Seg10, Sla10, Tur08, Wea87, Esp10b]. Reviews [HDS^{+}98, And10]. revolution [Kle05, Ano05, Kle05]. Ricerca [Maj29a]. ricerche [Maj33a]. Ricordo [Ama68b, Rec99c]. Riemann [GR87]. rifugio [Meo13]. righe [GM28]. Rigor [And10]. rinnovamento [DE06a, DE15]. Ritrovamento [Esp05b, Esp04]. ritrovato [Mau09]. Rivelazioni [Bel75]. rivista [GR12]. Robinson [HDS^{+}98]. Robotti [Bon09, RS09]. robust [SS12]. Roentgen [GM28, GMA06]. Rome [Seg88b]. rotante [GM28]. rotating [GM28]. rotation [GMA06]. Rowan [HDS^{+}98]. Rowan-Robinson [HDS^{+}98]. RuCl [BYK^{+}17]. Run [zXZ06]. Russian [Dir03]. Rydberg [GR06, Maj29a].

Sacha [Eis88, Wea87]. sagt [HDS^{+}98]. Salvatore [Mag16]. sarà [Mau09]. scalar [Esp07d, PW34]. Scattering [Maj06, BYK^{+}17, Gar84]. School [Zic66, Seg88b]. Sciascia [Eis88, HDS^{+}98, Wea87, Far07, Fra10a, Fra12, Nas07, RRR07]. science [Ano13, Car20, Far07, LL04, Nas07, LL04]. sciences [Fid02, Maj42, MM06b, Man05, Rec07b, Man06a]. Scientific [Bt06, Bt10, Bon09, Car20, Dir03, Fio08, Gue21, Rec06b, Rec07a, RE07, RS09, Ama68a, GR08a, Rec98, Rec99a, Rec99b, Rec99c, Rec06a]. scientifica [Ama68a, Rec99b, Rec99c]. Scientifiche [Fio08]. scientist [Ama66a, Bel75]. Scientists [Ste02]. scienza [Nas07]. Scienze [Rec07b, Fid02, Maj42]. scienziato [Bel75]. Scobel [HDS^{+}98]. scomparsa [Bel75, Bon10, Rec75a, Rec99c, Sci75, Sci77a, Sci82, Sci97, Sci01, Sci06]. scomparso [Bon02, Ron13]. Scope [Car20]. Scuola [Bon09]. sdoppiamento [GM28]. Search [GR06, Maj29a]. Searching [Esp12]. second [Bon04, Bon05, Bon10]. seconda [Bon10]. Secret [Cas13, Zie00]. Seems [GLR14]. Segrè [Lan17]. segunda [Bon05]. seinem [GHH02]. Selection [ABP98, APB96]. Self [Mar75, Rei12]. self-annihilating [Rei12]. Self-Consistent [Mar75]. semiconductor [Ali10, MZF^{+}12, Rei12, RLF12]. seminars [Rec78b]. Senatore [CGRS07]. sept [Ano05, Kle05]. September [Rec78b]. septième [CCJ^{+}34]. sessant’anni [Rec99c]. session [Rec78b]. seven [Kle05]. seven-times [Kle05]. Seventh [CCJ^{+}34]. seventy [Esp08]. shadows [Led01]. shifts [BS93]. SIF [Fio08]. Sighting [Sta17]. signature [DRM^{+}12, RLF12]. Signatures [MZF^{+}12]. simmetrica [Maj37]. Simplifying [vCK17]. Sinai [HDS^{+}98]. Singapore [Car20, Gue21]. singularities [Dir31]. sixty [Rec98, Rec99a, Rec99c]. skalaren [PW34]. slashers [Sci77a, Sci78]. Sobranie [Dir03]. Social [Rec07b, Fid02, Maj42, MM06b, Man05]. sociali [Fid02, Maj42, Rec07b]. solid [EF15, Rei12]. solid-state [EF15]. Solution [Sim06, Esp02]. Solutions [EN12]. Solvay [CCJ^{+}34, CCJ^{+}34]. solved [Kit06]. Some [ES10, Fer27, Rek83, Fer28, Spr91]. Sort [Ser12]. Source [Mag35]. sous [CCJ^{+}34]. space [Cha00, Heß94]. Spaceship [Sut92]. Spanish [Bon05, Sci78]. sparisce [Mau09]. specialist [Car20]. Spectra [Di 06]. speech [Ama66a]. Speed [DMA07]. sperimentale [Dra08]. spin [BYK^{+}17, Ing06]. spin-flip [Ing06]. Spinors [KT90, GR87, Heß94, Kal00]. splitting [DRM^{+}12, GMA06]. spontaneous [Gue13]. stability [Spr91]. standard [Tol10]. stars [Spr91]. state [EF15]. States [DE08]. Stat’i [Dir03, Gia83]. Stationary [DE08]. Statistical [DE04a, Fer27, Rec07b, Fer28, Fid02, GR05, Gue13, Maj42, MM06b, Man05, Man06a]. Statistiche [Rec07b, Fid02, Maj42]. statistico [Fer27]. statistische [Fer28]. steamer [GR10]. step [Rei12]. storica [Fin02]. story [Bri06, Esp07d]. stranger [Fra12]. Strong [Zic66]. Structure [CCJ^{+}34, Boh21, Hei34, CCJ^{+}34]. Stückelberg [DN05]. student [CGRS07]. studi [Nas07]. studio [GR12]. Study [BD02, Zic66, Gin13]. sua [Rec99c]. sul [GR10]. Sulla [Maj29b, Maj31c, Bel75, GR12, Maj33a, Rec75a]. Sulle [Pon87, GM28]. Sullo [GM28]. Superconducting [FK08]. superconductor [DRM^{+}12, MZF^{+}12, RLF12, SS12]. Superconductors [VOB09]. Superiore [Bon09]. sure [Hei34]. Surface [FK08]. surfaces [GR87, Kal00]. survey [Fin02, GHH02]. Swedish [Sci77b]. symmetric [MM06a]. Symmetrical [Maj37]. symmetry [Gue13]. system [Fer28]. Systems [Fer28, Gue13].

T [Dir03]. tachyons [Rec78b, Rec78b]. tassello [Ron11]. teacher [Car20, Esp06a]. Teaching [DE06b, DE07a, DE06a, DE15]. Teilchen [KK02]. Temperature [VOB09, Ing06]. Tenth [Man06b, Man06c]. tenu [CCJ^{+}34]. Teoria [Maj31d, Maj32c, Maj37, Maj28, Maj33a]. Teorica [Esp05b, DE06a, DE15, Dra08, Esp04, ME06]. Teoriya [Dir03]. termini [GM28, Maj31a]. terms [GM28, GMA06, Maj31a, MIM06]. Terrorism [Eis88]. testimonianza [CGRS07]. testimonianze [Rec87, RA91, RA00, RA02, Rec08, Rec11]. testimonies [Rec20, Gue21, Car20]. tests [Ano08]. Theoretical [DE06b, DE06c, DE07b, EMvdMR03, Esp04, Esp05b, Esp06d, Mag16, PR07, DE06a, DE07a, DE15, Dra06, Dra08, Esp06c, ERvdMB09, EAW15, Hei34, ME06]. Theorie [Fer28]. Theories [Bri65b]. théoriques [Hei34]. Theory [Dir03, Dvo97, Esp07b, Esp07c, GR09, Maj65, MA06, VOB09, Bia06, Esp09b, Fer28, Maj28, Maj32c, Maj33b, Maj33a, Maj37, MR06, MC06, MM06a, Spr91, Var02, Maj31d]. there [Fra04]. thin [Mau09]. Thomas [EFGP99, Esp02, Esp05a, GR05, GR08b, Mar75, Som32, Spr91]. time [Cha00]. times [ES10, Heß94, Kle05]. tolerant [DD13]. Tomography [HDS^{+}98]. topics [Rec78b]. Topological [FK08, DRM^{+}12]. Torino [Tur08]. tormented [Mag11]. Tra [Led01, Dra08, Nas07, RRR07]. Tracce [Pon87]. track [Pon87]. Trans [Eis88]. transformation [Wea76]. transforms [BK90]. transitions [BS93]. Translated [Wea87]. translation [Man05]. Transversale [Bon13]. transverse [Bon13]. treatment [DN05]. trigonal [DGS09]. triplet [Bia06]. triplets [Maj31d, MA06]. tripletti [Maj31d]. Troubled [And10, Clo10, Dom09, Esp10b, Nav11, Seg10, Sla10, Mag09]. Trudov [Dir03]. truth [Cas06c]. tunable [Ali10]. Tunis [Ron11]. Tunisi [Ron11]. Two [And10, EN12, Pav70]. Two-Electron [EN12]. Type [Ser12, Che12].

U [Mag16, HDS^{+}98]. Ugo [ACM11]. ultima [Cas06c]. ultimate [Cas06c]. ultra [Ing06]. ultra-low [Ing06]. undergraduate [Car20]. un’espressione [Maj29a]. unfinished [Cas13]. un’indagine [Fin02]. Unitary [Dir45, Wig39, Wig89]. Università [RCC^{+}06a, Maj87]. University [CGRS07, Maj87]. Unknown [Esp07b, Esp07c, Esp07d, ES10]. unoriented [Kal00]. Unpublished [RE07, Seg88a, Esp06g, ERvdMB09, Rec99b, Rec99c, Gen88]. Upset [Sul87]. useful [Che12].

vacant [Esp09a]. vacante [Esp09a]. Valeria [Tur08]. valevole [Maj29a]. valid [GR06, Maj29a]. valore [Fid02, Maj42, Rec07b]. value [Fid02, Maj42, MM06b, Man05, Man06a, Rec07b]. variabile [Maj32a, Maj32b]. Variable [BR45, Maj32a, Maj32b, MI06]. Variations [Esp06d, Esp06c]. verità [Cas06c]. Verschwinden [HDS^{+}98, Sci96, SW11, GHH02]. Via [Tur08, Sim06, Mal10]. vicenda [GR12]. vida [Mag11]. Viewpoint [Fra10b]. viii [RS09]. visit [Fer76]. vita [AL66, AL06, Rec09, Tur08]. vs [Ano08].

Walker [RS09]. warping [DGS09]. Wave [Cas06b, Cas06a, PW34]. wavefunction [RE98]. weak [Zic66]. Weinberg [Dvo97]. Weisskopf [Esp07d]. Wellengleichung [PW34]. Weyl [DE04b, KT90]. which [Fra12]. who [Bon02]. Wigner [Bri65b]. will [Fid02]. wires [Fra13]. Woolfson [HDS^{+}98]. Work [Rec06b, Ama68a, Rec99b, Rec99c, RE07, Rec09]. Works [Dir03]. World [Car20, Gue21]. wrapped [Cha00]. writing [Fra12]. Writings [ABP98, APB96].

X [HDS^{+}98]. X-ray [HDS^{+}98]. xii [RS09]. xxx [Wic81].

Years [Rog13, Esp08, GHH02, Mai06, Rec98, Rec99a, Rec99c]. York [Lan17, RS09, Seg10, Wea87]. young [Kle13].

Zener [DN05]. Zero [Ano17, DRM^{+}12, DGS09]. Zero-bias [DRM^{+}12]. zur [Fer28].

References

Amaldi:1998:CPE

[ABP98] Edoardo Amaldi, Giovanni Battimelli, and Giovanni Paoloni, editors. 20th Century Physics: Essays and Recollections: a Selection of Historical Writings, volume 3 of Edoardo Amaldi Foundation series. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1998. ISBN 981-02-2369-2. xiii + 747 pp. LCCN Q127.I8 A53 1998. URL http://www.

worldscientific.com/worldscibooks/10.1142/2852.

Arimondo:2010:EMB

[ACM10a] E. Arimondo, C. W. Clark, and W. C. Martin. Ettore Majorana and the birth of autoionization. In APS Meeting Abstracts: APS April Meeting 2010, February 13–16, 2010, page 10002. American Physical Society, Ridge, NY 11961, USA, February 2010. URL

http://adsabs.harvard.edu/abs/2010APS..APRH10002A. Ab-

stract #H10.002.

Arimondo:2010:CEM

[ACM10b] E. Arimondo, Charles W. Clark, and W. C. Martin. Colloquium: Ettore Majorana and the birth of autoionization. Reviews of Modern Physics, 82(3):1947–1958, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://adsabs.harvard.edu/abs/2010RvMP...82.1947A;
http://link.aps.org/doi/10.1103/RevModPhys.82.1947;
http://rmp.aps.org/abstract/RMP/v82/i3/p1947_1.

Arimondo:2011:EMU

[ACM11] Ennio Arimondo, Charles W. Clark, and William C. Martin. Ettore Majorana, Ugo Fano and autoionization. In APS Meeting Abstracts, page 24002. American Physical Society, Ridge, NY

11961, USA, March 2011. URL http://adsabs.harvard.edu/
abs/2011APS..MARV24002A.

Avignone:2008:DBD

[AEE08] Frank T. Avignone III, Steven R. Elliott, and Jonathan Engel. Double beta decay, Majorana neutrinos, and neutrino mass. Reviews of Modern Physics, 80(2):481–516, April 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.80.481;   http://rmp.aps.org/abstract/RMP/
v80/i2/p481_1.

Amaldi:1966:VOE

[AL66] Edoardo Amaldi and Rosario Liotta, editors. La vita e l’opera di Ettore Majorana: (1906–1938). Rariora. Accademia nazionale dei Lincei, Roma, Italia, 1966. xlix + 101 + 1 pp. LCCN ???? Reprinted in [AL06].

Amaldi:2006:VOE

[AL06] Edoardo Amaldi and Rosario Liotta, editors. La vita e l’opera di Ettore Majorana: (1906–1938), volume 1 of Rariora. Bardi, Roma, Italia, 2006. ISBN 88-88620-27-3. xlix + 101 + 1 pp. LCCN ???? Reprint of [AL66].

Alicea:2010:MFT

[Ali10] Jason Alicea. Majorana fermions in a tunable semiconductor device. Physical Review B: Condensed Matter and Materials Physics, 81(12):125318, March 2010. CODEN PRBMDO.

ISSN 1098-0121. URL http://link.aps.org/doi/10.1103/
PhysRevB.81.125318.

Alunni:2013:EMP

[Alu13] Charles Alunni. Ettore Majorana et la philosophie. (French) [Ettore Majorana and philosophy]. Revue de Synthèse, 134(1): 53–73, March 2013. ISSN 0035-1776 (print), 1955-2343 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11873-013-0215-x.

Amaldi:1966:EMM

[Ama66a] Edoardo Amaldi. Ettore Majorana, man and scientist: commemoration speech. In Zichichi [Zic66], pages 10–77. LCCN ????

Amaldi:1966:NBE

[Ama66b] Edoardo Amaldi. Nota biografica di Ettore Majorana. Accademia nazionale dei Lincei, Roma, Italy, 1966. xlix pp. LCCN ???? URL

http://books.google.com/books?id=WHq0AAAAIAAJ.

Amaldi:1968:LSE

[Ama68a] Edoardo Amaldi. L’opera scientifica di Ettore Majorana. (Italian) [The scientific work of Ettore Majorana]. Physis: Rivista Internazionale di Storia della Scienza, 10(??):173–187, 1968. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Amaldi::1968:REM

[Ama68b] Eduardo Amaldi. Ricordo di Ettore Majorana. (Italian) [I remember Ettore Majorana]. Storia della fisica, 9(??):300–318, ???? 1968.

Anderson:1932:AEE

[And32] Carl D. Anderson. The apparent existence of easily deflectable positives. Science, 76(1967):238–239, September 9, 1932.

CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/76/1967/

238. This article reports the first experimental discovery of the positron (positive electron).

Anderson:2010:TBM

[And10] Mark Anderson. Two beautiful minds, one perfectly rendered [reviews of “A Brillant Darkness: The Extraordinary Life and Mysterious Disappearance of Ettore Majorana, the Troubled Genius of the Nuclear Age (Magueijo, J.; 2009) and “Perfect Rigor: A Genius and the Mathematical Breakthrough of the Century” (Gessen, M.; 2009)]. IEEE Spectrum, 47(5):26, May 2010. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:EM

[Ano65] Anonymous. Ettore Majorana. Physics Today, 18(8):76, August 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://adsabs.harvard.edu/abs/1965PhT.
...18h..76..

Anonymous:1977:EM

[Ano77] Anonymous. Ettore Majorana. Bulletin Géodésique (Toulouse), 51(3):iii–iv, ???? 1977. CODEN BGDQAG. ISSN 0007-4632 (print), 1432-1394 (electronic). URL http://link.springer.

com/article/10.1007/BF02521602.

Anonymous:2005:CLB

[Ano05] Anonymous. Critique de livre: Il était sept Fois La révolution: Albert Einstein et les autres. La Recherche, ??(386):11–??, May 1, 2005. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-

9955 (electronic). URL http://www.larecherche.fr/idees/
livres/il-etait-sept-fois-revolution-albert-einstein-
autres-01-05-2005-73924.

Anonymous:2008:DVM

[Ano08] Anonymous. Dirac vs Majorana neutrinos: Low energy tests. AIP Conference Proceedings, 72:21–23, 2008. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic),

1935-0465. URL http://proceedings.aip.org/resource/2/
apcpcs/72/1/21_1.

Anonymous:2012:MFG

[Ano12] Anonymous. Majorana fermions: Glimpsed in a nanowire. Nature Nanotechnology, 8(??):??, ???? 2012. CODEN NNAABX.

ISSN 1748-3387 (print), 1748-3395 (electronic). URL http:

//www.nature.com/nnano/reshigh/2013/0113/full/nnano.
2012.252.html.

Anonymous:2013:EML

[Ano13] Anonymous. Ettore Majorana: de la légende à la science. (French) [Ettore Majorana: on the legend of science], volume 134(1) of Revue de synthèse: revue trimestrielle. Springer, Paris, France, 2013. ISBN 2-8178-0440-6. ISSN 0035-1776 (print), 1955-2343 (electronic). LCCN ????

Anonymous:2017:NZE

[Ano17] Anonymous. Neutrons zero in on elusive magnetic Majorana fermion. R&D Magazine, ??(??):??, June 9, 2017.

URL https://www.rdmag.com/news/2017/06/neutrons-zero-
elusive-magnetic-majorana-fermion?.

Amaldi:1996:ERC

[APB96] Edoardo Amaldi, Giovanni Paoloni, and Giovanni Battimelli, editors. Essays and recollections on 20th century physics: a selection of historical writings, volume 2 of Edoardo Amaldi Foundation series. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1996. ISBN 981-02-2369-2. xiii + 747 pp. LCCN Q127.I8 A53 1996.

Bergmann:2002:AIB

[BD02] Peter Gabriel Bergmann and Venzo De Sabbata, editors. Advances in the Interplay between Quantum and Gravity Physics: Proceedings of the NATO Advances Study Institute on Advances in the Interplay between Quantum and Gravity Physics, held in Erice, Italy, 30 April–10 May, 2001, volume 60 of NATO science series. Series II, Mathematics, physics, and chemistry. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 1-4020-0592-X. viii + 554 pp. LCCN QC178 .A38 2002. URL http://www.loc.gov/

catdir/enhancements/fy0822/2002069423-d.html;   http:
//www.loc.gov/catdir/toc/fy031/2002069423.html.

Bella:1975:RSS

[Bel75] Salvo Bella. Rivelazioni sulla scomparsa di uno scienziato, Ettore Majorana. (Italian) [Revelations on the disappearance of a scientist, Ettore Majorana]. Italia letteraria, Milano, Italia, 1975. ISBN ???? 157 pp. LCCN QC774.M34 B44. L3000.

Bianconi:2006:MTI

[Bia06] Antonio Bianconi. From the Majorana theory of ‘incomplete P′ triplets’ to Feshbach resonances. In Rapisarda et al. [RCC^{+}06a],

page ?? LCCN ???? URL http://pos.sissa.it/archive/
conferences/037/001/EMC2006_001.pdf.

Basu:1990:CIT

[BK90] Debabrata Basu and T. K. Kar. A class of integral transforms associated with the Majorana representation of SL(2, C). Journal of Mathematical Physics, 31(8):2035–2039, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i8/
p2035_s1.

Bohr:1921:ASb

[Boh21] Niels Bohr. Atomic structure. Nature, 108(2711):208–209, October 13, 1921. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v108/n2711/pdf/108208b0.pdf.

Bonolis:2002:MGS

[Bon02] Luisa Bonolis. Majorana: il genio scomparso. (Italian) [Majorana, the Genius who Disappeared], volume 27 of I grandi della scienza. Le Scienze, Milano, Italia, 2002. 111 pp. LCCN ????

Bonells:2004:DDM

[Bon04] Jordi Bonells. La deuxiéeme disparition de Majorana. (French) [The second disappearance of Majorana]. L. Levi, Paris, France, 2004. ISBN 2-86746-366-1 (paperback). 179 pp. LCCN ????

Bonells:2005:SDM

[Bon05] Jordi Bonells. La segunda desaparición de Majorana. (Spanish) [The second disappearance of Majorana]. Funambulista, Madrid, Spain, 2005. ISBN 84-934532-9-3. 216 pp. LCCN ???? Translation to Spanish by Lluís Agustí of [Bon04].

Bonolis:2009:BRF

[Bon09] Luisa Bonolis. Book review: Francesco Guerra and Nadia Robotti, Ettore Majorana. Aspects of his Scientific and Academic Activity. Pisa: Edizioni Scuola Normale Superiore, 2008. 243 pp., ISBN 978-88-7642-331-4. Nuncius, 24(2):540–541, ????

2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00912.

Bonells:2010:SSM

[Bon10] Jordi Bonells. La seconda scomparsa di Majorana. (Italian) [The second disappearance of Majorana], volume 6 of Passi. Keller, Rovereto (TN), Italia, 2010. ISBN 88-89767-12-X. 156 pp. LCCN ???? Translation to Italian by Angela Lorenzini of [Bon04].

Bontems:2013:LTE

[Bon13] Vincent Bontems. L’épistémologie transversale d’Ettore Majorana. (French) [The transverse epistomology of Ettore Majorana]. Revue de Synthèse, 134(1):29–51, March 2013. ISSN 0035-1776 (print), 1955-2343 (electronic). URL http://link.springer.

com/article/10.1007/s11873-013-0209-8.

Borthwick:1993:EMF

[Bor93] David Borthwick. Euclidean Majorana fermions, fermionic integrals, and relative Pfaffians. Journal of Mathematical Physics, 34 (7):2691–2712, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i7/p2691_s1.

Bloch:1945:AVM

[BR45] F. Bloch and I. I. Rabi. Atoms in variable magnetic fields. Reviews of Modern Physics, 17(2–3):237–244, April 1945. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.17.237;   http://rmp.aps.org/abstract/RMP/
v17/i2-3/p237_1.

Brink:1965:NF

[Bri65a] David Maurice Brink. Nuclear Forces, volume 354 of Commonwealth and international library: Selected readings in physics. Pergamon Press, New York, NY, USA; Oxford, UK, 1965. ISBN 0-08-011034-7. viii + 232 pp. LCCN QC173 .B8513 1965.

Brink:1965:THW

[Bri65b] David Maurice Brink. The theories of Heisenberg, Wigner and Majorana. In Nuclear Forces [Bri65a], chapter 2, pages 13–25. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965.

Brink:2006:SME

[Bri06] D. Brink. The story of Majorana’s exchange force. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN ???? URL http://pos.sissa.

it/archive/conferences/037/002/EMC2006_002.pdf.

Brouwer:2012:EMF

[Bro12] Piet W. Brouwer. Enter the Majorana fermion. Science, 336 (6084):989–990, May 25, 2012. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/41584897.

Bauch:1993:FSC

[BS93] Andreas Bauch and Roland Schröder. Frequency shifts in a cesium atomic clock due to Majorana transitions. Annalen der Physik (1900), 505(5):421–449, 1993. ISSN 1521-3889.

Bassani:2006:EMS

[Bt06] Giuseppe Franco Bassani and the Council of the Italian Physical Society, editors. Ettore Majorana: Scientific Papers: On Occasion of the Centenary of his Birth. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF). xliii + 282 pp. LCCN QC776 .M228 2006. URL http://

catdir.loc.gov/catdir/enhancements/fy0825/2006936896-
d.html;   http://catdir.loc.gov/catdir/enhancements/
fy0825/2006936896-t.html;   http://www.springerlink.
com/openurl.asp?genre=book\%26isbn=978-3-540-48091-4.

Bassani:2010:EMS

[Bt10] Giuseppe Franco Bassani and the Council of the Italian Physical Society, editors. Ettore Majorana: Scientific Papers: On Occasion of the Centenary of his Birth. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-08004-9 (paperback). xliii + 320 pp. LCCN ????

Breit:1935:NME

[BW35] G. Breit and E. Wigner. Note on Majorana’s exchange energy. Physical Review (2), 48(11):918–919, December 1, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.48.918.2.

Banerjee:2017:NSP

[BYK^{+}17] Arnab Banerjee, Jiaqiang Yan, Johannes Knolle, Craig A. Bridges, Matthew B. Stone, Mark D. Lumsden, David G. Mandrus, David A. Tennant, Roderich Moessner, and Stephen E. Nagler. Neutron scattering in the proximate quantum spin liquid — RuCl_{3}. Science, 356(6342):1055–1059, 2017. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

science.sciencemag.org/content/356/6342/1055.

Cabibbo:2006:NO

[Cab06] Nicola Cabibbo. Neutrino oscillations. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN ???? URL http://pos.sissa.

it/archive/conferences/037/003/EMC2006_003.pdf.

Caruso:2020:BRM

[Car20] Francisco Caruso. Book review: The Majorana Case: letters, documents, testimonies, by Erasmo Recami, Singapore, World Scientific, 2020, 400 pp., £29.90 (paperback), ISBN 978-981-12-0769-3. Scope: book, general interest. Level: undergraduate, advanced undergraduate, postgraduate, early career research, historian of science, teacher, and specialist. Contemporary Physics, 61 (4):304–305, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casalbuoni:2006:MICb

[Cas06a] R. Casalbuoni. Majorana and the infinite component wave equations. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN

???? URL http://pos.sissa.it/archive/conferences/037/
004/EMC2006_004.pdf.

Casalbuoni:2006:MICa

[Cas06b] Roberto Casalbuoni. Majorana and the infinite component wave equations. arxiv.org, ??(??):1–15, October 23, 2006. URL http:

//arxiv.org/abs/hep-th/0610252.

Castellani:2006:MMU

[Cas06c] Leandro Castellani. Mistero Majorana: l’ultima verità. (Italian) [The Majorana Mystery: the ultimate truth], volume 18 of Il Diforàno. Clinamen, Firenze, Italia, 2006. ISBN 88-8410-092-5. 96 pp. LCCN QC774.M34 C37 2006. URL http://catdir.loc.

gov/catdir/toc/casalini05/06400361.pdf.

Castellana:2013:UNS

[Cas13] Mario Castellana. The unfinished notebook. The secret life of Ettore Majorana. Revue de Synthèse, 134(1):125–128, ???? 2013. ISSN 0035-1776 (print), 1955-2343 (electronic).

Cockcroft:1934:SPN

[CCJ^{+}34] J. Cockcroft, J. Chadwick, F. Joliot, J. Joliot, N. Bohr, G. Gamov, P. A. M. Dirac, and W. Heisenberg, editors. Structure et propriétés des noyaux atomiques. Rapports et discussions du septième conseil de physique tenu à Bruxelles du 22 au 29 octobre 1933 sous les auspices de l’institut international de physique Solvay. (French) [Structure and properties of atomic nuclei. Reports and discussions of the Seventh Meeting on Physics held in Brussels from 22 to 29 October 1933 under the auspices of the Solvay International Institute of Physics]. Gauthier-Villars et cie, Paris, France, 1934. LCCN ???? Publiés par la commission administrative de l’institut.

Cennamo:2007:EMN

[CGRS07] Felice Cennamo, Francesco Guerra, Nadia Robotti, and Gilda Senatore. Ettore Majorana a Napoli: la testimonianza dell’allieva Gilda Senatore. (Italian) [Ettore Majorana at the University of Naples according to his student Gilda Senatore]. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 44(1):185– 202 (2008), 2007. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Chadwick:1932:EN

[Cha32a] James Chadwick. The existence of a neutron. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 136(830):692–708, June 1, 1932. ISSN 0950-1207 (print), 2053-9150 (electronic). URL http:

//www.jstor.org/stable/95816. This paper reports the discov-

ery of the neutron, for which Chadwick received the Nobel Prize in Physics in 1935.

Chadwick:1932:PEN

[Cha32b] James Chadwick. Possible existence of a neutron. Nature, 129 (3252):312, February 27, 1932. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v129/n3252/pdf/129312a0.pdf.

Chamblin:2000:EMF

[Cha00] Andrew Chamblin. Existence of Majorana fermions for M-branes wrapped in space and time. Journal of Mathematical Physics, 41 (8):5510–5516, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheung:2012:MEE

[Che12] Rebecca Cheung. Matter & energy: Elusive type of fermion reported: Majorana particles could be useful in quantum computing. Science News (Washington, DC), 181(10):11, 2012. CODEN SCNEBK. ISSN 1943-0930.

Close:2010:BRB

[Clo10] Frank Close. Book review: A Brilliant Darkness: The Extraordinary Life and Mysterious Disappearance of Ettore Majorana, the Troubled Genius of the Nuclear Age. Nature, 463(7277):33, January 6, 2010. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Cohen-Tannoudji:2012:NDP

[CT12] Gilles Cohen-Tannoudji. Les neutrinos défient les physiciens. (French) [Neutrinos challenge physicists]. La Recherche, ??(466): 38, June 19, 2012. CODEN RCCHBV. ISSN 0029-5671 (print),

1625-9955 (electronic). URL http://www.larecherche.fr/
physique-du-xxie-siecle/neutrinos-defient-physiciens-
01-07-2012-91376.

DAmaro:1998:EMD

[D’A98] S. D’Amaro. Ettore Majorana. a day in March. Ponte, 54(7): 138–140, ???? 1998. ISSN 0032-423X.

Deng:2013:FTQ

[DD13] Dong-Ling Deng and Lu-Ming Duan. Fault-tolerant quantum random-number generator certified by Majorana fermions. Physical Review A (Atomic, Molecular, and Optical Physics), 88(1): 012323:1–012323:6, July 2013. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://

link.aps.org/doi/10.1103/PhysRevA.88.012323.

DeGregorio:2005:PNO

[De 05] Alberto De Gregorio. Il ‘protone neutro’. Ovvero della laboriosa esclusione degli elettroni dal nucleo. (Italian) [The ‘neutral proton’. Namely, the laborious exclusion of electrons from the nucleus]. arxiv.org, ??(??):1–15, March 30, 2005. URL http://

arxiv.org/abs/physics/0603261.

DiGrezia:2004:FMS

[DE04a] Elisabetta Di Grezia and Salvatore Esposito. Fermi, Majorana and the statistical model of atoms. Foundations of Physics, 34(9): 1431–1450, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/B:FOOP.0000044099.18859.1d.

Drago:2004:FWQ

[DE04b] Antonino Drago and Salvatore Esposito. Following Weyl on quantum mechanics: The contribution of Ettore Majorana. Foundations of Physics, 34(5):871–887, May 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/adsabs.harvard.edu/abs/2004FoPh...34..871D;   http:/
/link.springer.com/article/10.1023/B:FOOP.0000022190.
47446.f3.

DeGregorio:2006:EFE

[DE06a] Alberto De Gregorio and Salvatore Esposito. Enrico Fermi e Ettore Majorana: continuitá e rinnovamento nell’insegnamento della fisica teorica. (Italian) [Enrico Fermi, Ettore Majorana: continuity and renewal in the teaching of theoretical physics]. arxiv.org, ??(??):1–13, February 1, 2006. URL http://

adsabs.harvard.edu/abs/2006physics...2008D;   http://
arxiv.org/abs/physics/0602008. To be published in Sapere

[(Bari) (ISSN 0036-4681)].

DeGregorio:2006:TTP

[DE06b] Alberto De Gregorio and Salvatore Esposito. Teaching theoretical physics: the cases of Enrico Fermi and Ettore Majorana. arxiv.org, ??(??):21, February 21, 2006. URL http://arxiv.

org/abs/physics/0602146.

Drago:2006:LAM

[DE06c] Antonino Drago and Salvatore Esposito. A logical analysis of Majorana’s papers of theoretical physics. Electronic Journal of Theoretical Physics, 3(10):249–263, ???? 2006. CODEN EJTPBC.

ISSN 1729-5254. URL http://adsabs.harvard.edu/abs/
2006EJTP....3j.249D;   http://www.ejtp.info/articles/

ejtpv3i10p249.pdf. Special Majorana Centenary Issue: Majorana Legacy in Contemporary Physics.

DeGregorio:2007:TTP

[DE07a] Alberto De Gregorio and Salvatore Esposito. Teaching theoretical physics: The cases of Enrico Fermi and Ettore Majorana. American Journal of Physics, 75(9):781–790, September 2007. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Drago:2007:EMC

[DE07b] Antonino Drago and Salvatore Esposito. Ettore Majorana’s course on theoretical physics: a recent discovery. Physics in Perspective (PIP), 9(3):329–345, September 2007. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/s00016-006-
0305-3;   http://www.arxiv.org/abs/physics/0503084.

DiGrezia:2008:MQS

[DE08] Elisabetta Di Grezia and Salvatore Esposito. Majorana and the quasi-stationary states in nuclear physics. Foundations of Physics, 38(3):228–240, March 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

adsabs.harvard.edu/abs/2008FoPh...38..228D;   http://
link.springer.com/article/10.1007/s10701-007-9200-2.

DeGregorio:2015:EFE

[DE15] Alberto De Gregorio and Salvatore Esposito. Enrico Fermi e Ettore Majorana: continuità e rinnovamento nell’insegnamento della fisica teorica. (Italian) [Enrico Fermi and Ettore Majorana: continuity and renewal in the teaching of theoretical physics]. Report, L’Università di Roma ^{≪}La Sapienza^{≫}e l Università di Napoli^{≪}Federico II^{≫}, Roma, Italia e Napoli, Italia, July 27, 2015.

13 pp. URL https://www.researchgate.net/publication/
2175080.

Dora:2009:MZM

[DGS09] Balázs Dóra, Miklós Gulácsi, and Pasquale Sodano. Majorana zero modes in graphene with trigonal warping. Physica Status Solidi (RRL) Rapid Research Letters, 3(6):169–171, 2009. ISSN 1862-6254 (print), 1862-6270 (electronic).

DiGrezia:2006:MII

[Di 06] Elisabetta Di Grezia. Majorana and the investigation of infrared spectra of ammonia. Electronic Journal of Theoretical Physics, 3(10):225–238, ???? 2006. CODEN EJTPBC.

ISSN 1729-5254. URL http://adsabs.harvard.edu/abs/
2006EJTP....3j.225D;   http://www.ejtp.com/articles/

ejtpv3i10p225.pdf. Special Majorana Centenary Issue: Majorana Legacy in Contemporary Physics.

Dirac:1930:AEP

[Dir30] P. A. M. Dirac. On the annihilation of electrons and protons. Mathematical proceedings of the Cambridge Philosophical Society, 26(3):361–375, 1930. CODEN MPCPCO. ISSN 0305-0041, 1469-8064.

Dirac:1931:QSE

[Dir31] P. A. M. Dirac. Quantised singularities in the electromagnetic field. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 133(821):60–72, September 1, 1931. CODEN PRLAAZ. ISSN 0080-4630. URL http://www. jstor.org/stable/95639. Russian translation in [Dir03, Vol. II, pp. 388–??].

Dirac:1945:URL

[Dir45] P. A. M. Dirac. Unitary representations of the Lorentz group. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 183(994):284–295, February 22, 1945. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/97721.

Dirac:2003:SNT

[Dir03] P. A. M. Dirac. Sobranie Nauchnykh Trudov T. II. Kvantovaya Teoriya (Nauchnye Stat’i: 1924–1947). (Russian) [Collected Scientific Works. Quantum Theory (Scientific Papers: 1924–1947)]. Klassiki Nauki. Fizmatlit, Moscow, Russia, 2003. ISBN 5-9221-0381-4. 846 pp. LCCN ???? Edited by A. D. Sukhanov.

Dragoni:2007:QME

[DMA07] Giorgio Dragoni, Giulio Maltese, and Luisa Atti. Quirino Majorana’s experiments on the speed of light and gravitational absorption. Physics in Perspective (PIP), 9(3):281–304, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-006-0304-4.

DiGiacomo:2005:MFL

[DN05] Francesco Di Giacomo and Evgenii E. Nikitin. The Majorana formula and the Landau–Zener–Stückelberg treatment of the avoided crossing problem. Physics-Uspekhi, 48(5):515–517, May 2005. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780

(electronic). URL http://ufn.ru/en/articles/2005/5/e/.

Dominy:2009:BRB

[Dom09] Margaret F. Dominy. Book review: A Brilliant Darkness: The Extraordinary Life and Mysterious Disappearance of Ettore Majorana, the Troubled Genius of the Nuclear Age. Library Journal, 134(18):87, November 2009. CODEN LIBJA7. ISSN 0000-0027, 0363-0277.

Dragoni:2006:MGQ

[Dra06] Giorgio Dragoni. E. Majorana as a guide in Quirino Majorana’s experiments. Original letters and documents on an experimental and theoretical collaboration. In Rapisarda et al. [RCC^{+}06a],

page ?? LCCN ???? URL http://pos.sissa.it/archive/
conferences/037/005/EMC2006_005.pdf.

Dragoni:2008:EQM

[Dra08] Giorgio Dragoni. Ettore e Quirino Majorana: tra fisica teorica e sperimentale. (Italian) [Ettore and Quirino Majorana: between theoretical and experimental physics]. Consiglio Nazionale delle Ricerche, Roma, Italia, 2008. ISBN 88-7438-044-5. l + 368 pp. LCCN QC16.M335. 50.00 EUR.

URL http://www.gbv.de/dms/casalini/09/09923732.pdf;
http://www.sif.it/libri/dragoni.

Das:2012:ZBP

[DRM^{+}12] Anindya Das, Yuval Ronen, Yonatan Most, Yuval Oreg, Moty Heiblum, Hadas Shtrikman, et al. Zero-bias peaks and splitting in an Al–InAs nanowire topological superconductor as a signature of Majorana fermions. Nature Physics, 8(12):887–895, November 11, 2012. CODEN NPAHAX. ISSN 1745-2473 (print),

1745-2481 (electronic). URL http://www.nature.com/nphys/
journal/v8/n12/full/nphys2479.html.

DeGregorio:2009:EFO

[DS09] Alberto De Gregorio and Fabio Sebastiani. Enrico Fermi and the old quantum physics. arxiv.org, pages 1–12, April 1, 2009. URL

http://arxiv.org/abs/0904.0260.

DeGregorio:2010:EFO

[DS10] A. De Gregorio and F. Sebastiani. Enrico Fermi and the old quantum physics. Quaderni di Storia della fisica, 16:81–102, 2010. CODEN ???? ISSN 1594-9974 (print), 1827-6164 (electronic).

URL http://en.sif.it/journals/qsf/econtents.

Dunne:1996:MP

[Dun96] Robert S. Dunne. Majorana parafermions. Journal of Mathematical Physics, 37(1):452–459, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dvoeglazov::1997:AMI

[Dvo97] Valeri V. Dvoeglazov. Another Majorana idea — real and imaginary in the Weinberg theory. International Journal of Theoretical Physics, 36(3):635–642, March 1997. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic). URL http://link.

springer.com/article/10.1007/BF02435883.

Esposito:2015:PEM

[EAW15] Salvatore Esposito, E. (Evgeny) Akhmedov, and Frank Wilczek. The physics of Ettore Majorana: phenomenological, theoretical, and mathematical. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-04402-2 (hardcover), 1-316-19108-7 (PDF ebook). xi + 382 pp. LCCN QC19.6 .E87 2015.

Elliott:2015:CMF

[EF15] Steven R. Elliott and Marcel Franz. Colloquium: Majorana fermions in nuclear, particle, and solid-state physics. Reviews of Modern Physics, 87(1):137–??, January/March 2015. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.137.

Epele:1999:PAA

[EFGP99] L. N. Epele, H. Fanchiotti, C. A. García Canal, and J. A. Ponciano. Padé approximant approach to the Thomas–Fermi problem. Physical Review A (Atomic, Molecular, and Optical Physics), 60(1):280–283, July 1999. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http:

//link.aps.org/doi/10.1103/PhysRevA.60.280.

Eisenzweig:1988:TLR

[Eis88] Uri Eisenzweig. Terrorism in life and in real literature [book review: Leonardo Sciascia, The Moro Affair and The Mystery of Majorana. Trans. Sacha Rabinovitch]. Diacritics, 18(3):32–42, Autumn 1988. ISSN 0300-7162 (print), 1080-6539 (electronic).

URL http://www.jstor.org/stable/465253.

Esposito:2003:EMN

[EMvdMR03] Salvatore Esposito, Ettore Majorana, Jr., Alwyn van der Merwe, and Erasmo Recami, editors. Ettore Majorana: Notes on Theoretical Physics, volume 133 of Fundamental Theories of Physics. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1649-2, 90-481-6435-4 (print), 94-017-0107-5 (e-book). xxv + 486 pp.

LCCN QC19.2-20.85. URL http://link.springer.com/book/
10.1007/978-94-017-0107-5/page/1.

Esposito:2012:MST

[EN12] Salvatore Esposito and Adele Naddeo. Majorana solutions to the two-electron problem. Foundations of Physics, 42 (12):1586–1608, December 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

adsabs.harvard.edu/abs/2012FoPh...42.1586E;   http://
link.springer.com/article/10.1007/s10701-012-9685-1.

Esposito:2009:EMU

[ERvdMB09] Salvatore Esposito, Erasmo Recami, Alwyn van der Merwe, and Roberto Battiston, editors. Ettore Majorana: unpublished research notes on theoretical physics, volume 159 of Fundamental theories of physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9113-3, 1-4020-9114-1 (e-book). lvi + 451 pp. LCCN QC3 .M273 2009.

URL http://link.springer.com/book/10.1007/978-1-4020-
9114-8/page/1.

Esposito:2010:FTL

[ES10] Salvatore Esposito and Giovanni Salesi. Fundamental times, lengths and physical constants: Some unknown contributions by Ettore Majorana. Annalen der Physik (1900), 522(7):456–466, July 2010. ISSN 0003-3804 (print), 1521-3889 (electronic). URL

http://adsabs.harvard.edu/abs/2010AnP...522..456E.

Esposito:1998:CMF

[Esp98] Salvatore Esposito. Covariant Majorana formulation of electrodynamics. Foundations of Physics, 28(2):231–244, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A:1018752803368.

Esposito:2002:MST

[Esp02] Salvatore Esposito. Majorana solution of the Thomas–Fermi equation. American Journal of Physics, 70(8):852–856, August 2002. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v70/
i8/p852_s1.

Esposito:2004:CFT

[Esp04] S. Esposito. Il corso di fisica teorica di Ettore Majorana: il ritrovamento del Documento Moreno. (Italian) [The course of theoretical physics of Ettore Majorana: the discovery of the Moreno Paper]. arxiv.org, ??(??):1–18, November

2004. URL http://adsabs.harvard.edu/abs/2004physics.
.11199E;   http://arxiv.org/abs/physics/0411199. Pub-

lished in [Esp05b].

Esposito:2005:AMT

[Esp05a] Salvatore Esposito. Again on Majorana and the Thomas–Fermi model: a comment to physics/0511222. arxiv.org, ??(??):1–

5, December 5, 2005. URL http://adsabs.harvard.edu/abs/
2005physics..12259E;   http://arxiv.org/abs/physics/

0512259. See [GR05].

Esposito:2005:CFT

[Esp05b] Salvatore Esposito. Il corso di fisica teorica di Ettore Majorana: il ritrovamento del documento Moreno. (Italian) [Ettore Majorana’s course on theoretical physics: The finding of the Moreno paper]. Il Nuovo Saggiatore, 21(??):21–33, ???? 2005. ISSN 0393-4578 (print), 1827-6148 (electronic). URL http://prometeo.

sif.it:8080/papers/online/sag/021/03-04/pdf/04.pdf.

Esposito:2006:MRT

[Esp06a] Salvatore Esposito. E. Majorana researcher and teacher: the latest achievements. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN

???? URL http://pos.sissa.it/archive/conferences/037/
006/EMC2006_006.pdf.

Esposito:2006:EMF

[Esp06b] Salvatore Esposito. Ettore Majorana — fleeting genius. Physics World, 19(8):34–36, August 2006. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

Esposito:2006:FVTa

[Esp06c] Salvatore Esposito. Four variations on theoretical physics by Ettore Majorana. arxiv.org, ??(??):1–23, April 7, 2006. URL http:

//arxiv.org/abs/physics/0604064.

Esposito:2006:FVTb

[Esp06d] Salvatore Esposito. Four variations on theoretical physics by Ettore Majorana. Electronic Journal of Theoretical Physics, 3(10): 265–283, ???? 2006. CODEN EJTPBC. ISSN 1729-5254. URL

http://adsabs.harvard.edu/abs/2006EJTP....3j.265E;
http://www.ejtp.com/articles/ejtpv3i10p265.pdf. Special

Majorana Centenary Issue: Majorana Legacy in Contemporary Physics.

Esposito:2006:MPI

[Esp06e] Salvatore Esposito. Majorana and the path-integral approach to quantum mechanics. arxiv.org, ??(??):1–9, March 17,

2006. URL http://adsabs.harvard.edu/abs/2006physics..
.3140E;   http://arxiv.org/abs/physics/0603140.

Esposito:2006:PLE

[Esp06f] Salvatore Esposito. A peculiar lecture by Ettore Majorana. European Journal of Physics, 27(5):1147–1156, 2006. CO-DEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2006EJPh...
27.1147E;   http://stacks.iop.org/0143-0807/27/i=5/a=
014.

Esposito:2006:MIF

[Esp06g] Salvatore Esposito. Un manoscritto inedito in francese di Ettore Majorana. (Italian) [An unpublished manuscript in French by Ettore Majorana]. arxiv.org, ??(??):1–16, July

2006. URL http://adsabs.harvard.edu/abs/2006physics..
.7099E;   http://arxiv.org/abs/physics/0607099.

Esposito:2006:LPE

[Esp06h] Salvatore Esposito. Una lezione particolare di Ettore Majorana. (Italian) [A peculiar lecture by Ettore Majorana]. arXiv.org,

pages 1–16, 2006. URL http://arxiv.org/abs/physics/
0512174.

Esposito:2007:EMC

[Esp07a] Salvatore Esposito. Ettore Majorana centennial and neutrino legacy. Nuclear Physics B, Proceedings Supplements, 168:377– 382, June 2007. CODEN NPBSE7. ISSN 0920-5632 (print),

1873-3832 (electronic). URL http://adsabs.harvard.edu/
abs/2007NuPhS.168..377E;   http://www.sciencedirect.
com/science/article/pii/S0920563207001855.

Esposito:2007:HTQa

[Esp07b] Salvatore Esposito. Hole theory and quantum electrodynamics in an unknown manuscript in French by Ettore Majorana. Foundations of Physics, 37(6):956–976, June 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

adsabs.harvard.edu/abs/2007FoPh...37..956E;   http://
link.springer.com/article/10.1007/s10701-007-9136-6.

Esposito:2007:HTQb

[Esp07c] Salvatore Esposito. Hole theory and quantum electrodynamics in an unknown manuscript in French by Ettore Majorana. Foundations of Physics, 37(7):1049–1068, July 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

adsabs.harvard.edu/abs/2007FoPh...37.1049E;   http://
link.springer.com/article/10.1007/s10701-007-9141-9.

Esposito:2007:USM

[Esp07d] Salvatore Esposito. An unknown story: Majorana and the Pauli– Weisskopf scalar electrodynamics. Annalen der Physik (1900), 16 (12):824–841, December 2007. ISSN 1521-3889.

Esposito:2008:EMH

[Esp08] Salvatore Esposito. Ettore Majorana and his heritage seventy years later. Annalen der Physik (1900), 17(5):302–318, May 2008. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://

adsabs.harvard.edu/abs/2008AnP...520..302E;   http://
onlinelibrary.wiley.com/doi/10.1002/andp.200810296/.

Esposito:2009:CVE

[Esp09a] Salvatore Esposito. La cattedra vacante: Ettore Majorana: ingegno e misteri. (Italian) [The vacant chair: Ettore Majorana: genius and mysteries]. Script. Liguori, Napoli, Italia, 2009. ISBN 88-207-4691-3. vii + 253 pp. LCCN QC774.M34. 17.50 EUR.

URL http://www.gbv.de/dms/casalini/09/0974407X.pdf.

Esposito:2009:TFE

[Esp09b] Salvatore Esposito. A theory of ferromagnetism by Ettore Majorana. Annals of Physics, 324(1):16–29, January 2009. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic). URL

http://adsabs.harvard.edu/abs/2009AnPhy.324...16E.

Esposito:2010:DEM

[Esp10a] Salvatore Esposito. The disappearance of Ettore Majorana: an analytic examination. Contemporary Physics, 51 (3):193–209, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic). URL http://

adsabs.harvard.edu/abs/2010ConPh..51..193E;   http://
www.tandfonline.com/doi/abs/10.1080/00107510903510196.

Esposito:2010:MMB

[Esp10b] Salvatore Esposito. The Majorana mystery [book review: A Brilliant Darkness: The Extraordinary Life and Mysterious Disappearance of Ettore Majorana, the Troubled Genius of the Nuclear Age, João Magueijo, 2009, Basic Books, £15.99/$27.50 (hb), 304pp]. Physics World, 23(3):44–45, March 2010. CO-DEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://physicsworldarchive.iop.org/full/
pwa-pdf/23/03/phwv23i03a42.pdf. See response [Rec10].

Esposito:2012:SED

[Esp12] Salvatore Esposito. Searching for an equation: Dirac, Majorana and the others. Annals of Physics, 327(6):1617–1644, June 2012. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0003491612000346.

Farrell:2007:ESL

[Far07] Joseph Farrell. The ethics of science: Leonardo Sciascia and the Majorana case: (Ettore Majorana). Modern Language Review, 102(??):1021–??, ???? 2007. ISSN 0026-7937 (print), 2222-4319

(electronic). URL http://www.jstor.org/stable/20467548.

Fermi:1927:MSP

[Fer27] Enrico Fermi. Un metodo statistico per la determinazione di alcune proprietá dell’atomo. (Italian) [A statistical method for the determination of some atomic properties]. Rendiconti dell’Accademia Nazionale dei Lincei, 6(6):602–607, ???? 1927. CODEN AANLAW. ISSN 0001-4435. This paper is the Fermi side of the Thomas–Fermi model, which Fermi developed independently a year after the work of Llewellyn H. Thomas ([Tho27]). See [Seg70, page 53], [GR08b], and [DS09] for historical background.

Fermi:1928:SMB

[Fer28] Enrico Fermi. Eine statistische Methode zur Bestimmung einiger Eigenschaften des Atoms und ihre Anwendung auf die Theorie des periodischen Systems der Elemente. (German) [A statistical method for the determination of some atomic properties and its application to the theory of the periodic system of elements]. Zeitschrift für Physik, 48(1–2):73–79, January 1928. CO-

DEN ZEPYAA. ISSN ???? URL http://www.springerlink.
com/content/v762582061464612/. See English translation in

[Mar75, pp. 205–213].

Fermi:1976:VEM

[Fer76] Laura Fermi. A visit to the “Ettore Majorana” Centre. Physics Today, 29(8):9–13, August 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

adsabs.harvard.edu/abs/1976PhT....29h...9F;   http://
www.physicstoday.org/resource/1/phtoad/v29/i8/p9_s1.

Fidomanzo:2002:VDL

[Fid02] Vincenzo Fidomanzo. “Il valore delle leggi statistiche nella fisica e nelle scienze sociali” di Ettore Majorana. Ovvero: una percezione del “liberto” arbitrio. (Italian) [“The value of statistical laws in physics and in the social sciences” of Ettore Majorana. that is, a perception of “free will”]. Revue européenne des sciences sociales = European Journal of Social Sciences, 40(122):155–238, ???? 2002. ISSN 0048-8046 (print), 1663-4446 (electronic). URL http:

//www.jstor.org/stable/i40016249.

Finzi:2002:EMU

[Fin02] Roberto Finzi. Ettore Majorana: un’indagine storica. (Italian) [Ettore Majorana: an historical survey], volume 86 of Letture di pensiero e d’arte. Edizioni di storia e letteratura, Roma, Italy, 2002. ISBN 88-8498-062-3. 114 pp. LCCN QC774.M34 F56 2002.

Fiorini:2006:NDM

[Fio06] Ettore Fiorini. Neutrinos: Dirac or Majorana? In Rapisarda et al. [RCC^{+}06a], page ?? LCCN ???? URL http://pos.

sissa.it/archive/conferences/037/007/EMC2006_007.pdf.

Fiorini:2007:MNM

[Fio07] Ettore Fiorini. Measurement of neutrino mass in double beta decay. Europhysics News, 38(4):30–34, July/August 2007. CO-DEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL http://www.europhysicsnews.org/articles/
epn/abs/2007/04/epn07404/epn07404.html.

Fiorini:2008:BRB

[Fio08] Ettore Fiorini. Book review: Ettore Majorana, Scientific Papers, F. Bassani Giuseppe, SIF Edizioni Scientifiche, 2006, 320 pp., ISBN 88-7438-031-3. Europhysics News, 39(2):31, March/ April 2008. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-

1092 (electronic). URL http://www.europhysicsnews.org/
articles/epn/pdf/2008/02/Whole_issue.pdf.

Fu:2008:SPE

[FK08] Liang Fu and C. L. Kane. Superconducting proximity effect and Majorana fermions at the surface of a topological insulator. Physical Review Letters, 100(9):096407, March 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://link.aps.org/doi/10.1103/PhysRevLett.
100.096407.

Fradkin:1966:CPM

[Fra66] D. M. Fradkin. Comments on a paper by Majorana concerning elementary particles. American Journal of Physics, 34(4):314–318, April 1966. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://ajp.aapt.org/resource/1/ajpias/

v34/i4/p314_s1. Reprinted in [Fra06]. Esposito [Esp08] reports about this paper: “The importance of Majorana’s article [Maj32c] was first realized by B. L. van der Waerden but, unfortunately, it remained unnoticed for more than three decades until D. Fradkin, informed by Amaldi, reexamined that pioneering work in light of later developments and clearly explained the relevance of Majorana’s approach and results accomplished many years earlier”.

Franklin:2004:TRN

[Fra04] Allan Franklin. Are there really neutrinos?: an evidential history. Perseus Publishers, Cambridge, MA, USA, 2004. ISBN 0-7382-0265-7, 0-8133-4128-0 (paperback). ix + 371 pp. LCCN QC793.5.N42 F73 2004. URL http://www.loc.gov/

catdir/enhancements/fy0837/2008530710-d.html;   http:
//www.loc.gov/catdir/toc/fy0805/2008530710.html.

Fradkin:2006:CPM

[Fra06] D. M. Fradkin. Comments on a paper by Majorana concerning elementary particles. Electronic Journal of Theoretical Physics, 3(10):305–314, ???? 2006. CODEN EJTPBC.

ISSN 1729-5254. URL http://adsabs.harvard.edu/abs/
2006EJTP....3j.305F;   http://www.ejtp.com/articles/

ejtpv3i10p305.pdf. Special Majorana Centenary Issue: Majorana Legacy in Contemporary Physics. Reprint of [Fra66].

Francese:2010:LSB

[Fra10a] Joseph Francese. Leonardo Sciascia and The Disappearance of Majorana. Journal of Modern Italian Studies, 15 (5):715–733, 2010. ISSN 1354-571X (print), 1469-9583 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/1354571X.2010.523541.

Franz:2010:VRM

[Fra10b] Marcel Franz. Viewpoint: Race for Majorana fermions. Physics, 3(??):24–26, March 15, 2010. URL http://link.aps.org/doi/

10.1103/Physics.3.24;   http://physics.aps.org/articles/
v3/24.

Francese:2012:LSB

[Fra12] Joseph Francese. Leonardo Sciascia’s L’affaire Moro: Re-writing fact, which can be stranger than fiction. Modern Italy, 17(3):383– 398, ???? 2012. ISSN 1353-2944 (print), 1469-9877 (electronic).

URL http://www.tandfonline.com/loi/cmit20.

Franz:2013:MW

[Fra13] Marcel Franz. Majorana’s wires. Nature Nanotechnology, 8(3): 149–152, March 5, 2013. CODEN NNAABX. ISSN 1748-3387 (print), 1748-3395 (electronic). URL http://www.nature.com/

nnano/journal/v8/n3/full/nnano.2013.33.html.

Garavaglia:1984:DMN

[Gar84] Theodore Garavaglia. Dirac- and Majorana-neutrino-mass effects in neutrino-electron elastic scattering. Physical Review D (Particles and Fields), 29(3):387–392, February 1984. CODEN PRV-DAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Gentile:1988:LIE

[Gen88] B. Gentile. Lettere inedite di Ettore Majorana a Giovanni Gentile, Jr.. (Italian) [Unpublished correspondence of Ettore Majorana to Giovanni Gentile]. Giornale Critico della Filosofia Italiana, 8(2):145–153, August 1988. ISSN 0017-0089.

Goldhaber:2011:CDM

[GG11] Alfred Scharff Goldhaber and Maurice Goldhaber. Clarifying Dirac and Majorana distinctions. Physics Today, 64(7):12, July 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/
1/phtoad/v64/i7/p12_s3.

Giacomuzzi:2002:EMB

[GHH02] Salvatore Matteo Giacomuzzi, Gerhard Holzmüller, and Gerhard Huemer. Ettore Majorana (1906–1938): Eine Bestandsaufnahme 64 Jahre nach seinem Verschwinden. (German) [Ettore Majorana (1906–1938): A survey 64 years after his disappearance]. Berichte zur Wissenschaftsgeschichte, 25(2):137–148, June 2002. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Gianetto:1983:EMP

[Gia83] Enrico Gianetto. Ettore Majorana e il problema degli stati ad energia negativa. Giornale di Fisica, 34(2):151–158, ???? 1983. CO-DEN GFSIAD. ISSN 0017-0283 (print), 1827-6156 (electronic).

Gianetto:1985:MOF

[Gia85] Enrico Gianetto. A Majorana–Oppenheimer formulation of quantum electrodynamics. Lettere al Nuovo Cimento Series 2, 44(3): 140–144, October 1, 1985. CODEN LNUCAE. ISSN 0375-930X (print), 1827-613X (electronic). URL http://link.springer.

com/article/10.1007/BF02746912.

Gingras:2013:CPE

[Gin13] Yves Gingras. La carrière des publications d’Ettore Majorana: une Étude bibliométrique. (French) [The publication career of Ettore Majorana: a bibliometric study]. Revue de Synthèse, 134 (1):75–87, March 2013. ISSN 0035-1776 (print), 1955-2343 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11873-013-0210-2.

Guerra:2014:WEC

[GLR14] Francesco Guerra, Matteo Leone, and Nadia Robotti. When energy conservation seems to fail: The prediction of the neutrino. Science & Education (Springer), 23(6):1339–1359, June 2014. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Gentile:1928:SST

[GM28] Giovannino Gentile, Jr. and Ettore Majorana. Sullo sdoppiamento dei termini Roentgen e ottici a causa dell’elettrone rotante e sulle intensità delle righe del cesio. (Italian) [On the doubling of Roentgen and optical terms due to the rotating electron, and on the intensities of cesium lines]. Atti della Reale Accademia Nazionale dei Lincei. Rendiconti Classe di Scienze fisiche, matematiche e naturali, 8(??):229–233, ???? 1928. Presented by Orso Mario Corbino at the meeting of the Academy on July 24, 1928.

Gentile:2006:SRO

[GMA06] G. Gentile, E. Majorana, and Ennio Arimondo. On the splitting of the Roentgen and optical terms caused by the electron rotation and on the intensity of the cesium lines. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF). LCCN QC776 .M228

2006. URL http://link.springer.com/content/pdf/10.
1007/978-3-540-48095-2_1.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-48095-2/1/1.pdf. Scientific

paper number 1a.

Goodstein:1999:CHB

[Goo99] Judith Goodstein. A conversation with Hans Bethe. Physics in Perspective (PIP), 1(3):253–281, October 1999. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050022;   http://www.springerlink.com/content/
k4f0pllveakla47q/.

Grinstein:1987:DMS

[GR87] Benjamin Grinstein and Ryan Rohm. Dirac and Majorana spinors on nonorientable Riemann surfaces. Communications in Mathematical Physics, 111(4):667–675, ???? 1987. CODEN CM-PHAY. ISSN 0010-3616 (print), 1432-0916 (electronic). URL

http://projecteuclid.org/euclid.cmp/1104159727.

Guerra:2005:FPE

[GR05] Francesco Guerra and Nadia Robotti. A forgotten publication of Ettore Majorana on the improvement of the Thomas– Fermi statistical model. arxiv.org, ??(??):1–17, November 28,

2005. URL http://adsabs.harvard.edu/abs/2005physics..
11222G;   http://arxiv.org/abs/physics/0511222. See re-

buttal in [Esp05a].

Guerra:2006:MDE

[GR06] Francesco Guerra and Nadia Robotti. Majorana dr Ettore: Search for a general expression of Rydberg corrections, valid for neutral atoms or positive ions. In Bassani and the Council of the Italian Physical Society [Bt06], pages 21–36. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF).

LCCN QC776 .M228 2006. URL http://link.springer.com/
content/pdf/10.1007/978-3-540-48095-2_2.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-540-48095-2/

2/1.pdf. Scientific paper number 1b. Reprint of the Italian original [Maj29a], with an English translation.

Guerra:2008:EMA

[GR08a] Francesco Guerra and Nadia Robotti. Ettore Majorana: aspects of his scientific and academic activity, volume 6 of CRM Series. Edizioni della Normale, Pisa, Italia, 2008. ISBN 88-7642-331-1 (paperback), 88-7642-331-4 (paperback). ISSN 2239-4524. xii + 243 pp. LCCN ???? URL http://www.maa.

org/publications/maa-reviews/ettore-majorana-aspects-
of-his-scientific-and-academic-activity.

Guerra:2008:EMF

[GR08b] Francesco Guerra and Nadia Robotti. Ettore Majorana’s forgotten publication on the Thomas–Fermi model. Physics in Perspective (PIP), 10(1):56–76, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:/

/link.springer.com/article/10.1007/s00016-007-0340-8;

 http://www.springerlink.com/content/53l26g8353l2421g/

.

Guerra:2009:EFD

[GR09] Francesco Guerra and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: The influence of his theory of beta decay. Physics in Perspective (PIP), 11(4): 379–404, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://adsabs.harvard.

edu/abs/2009PhP....11..379G;   http://link.springer.
com/article/10.1007/s00016-008-0415-1;   http://www.
springerlink.com/content/65w213132524072n/.

Guerra:2010:EMS

[GR10] Francesco Guerra and Nadia Robotti. Ettore Majorana sul piroscafo. (Italian) [Ettore Majorana on the steamer]. Il Nuovo Saggiatore, 26(5–6):??, ???? 2010. ISSN 0393-4578 (print), 1827-6148 (electronic). URL http://static.sif.

it:8080/SIF/resources/public/files/opinioni/op_1011_
guerra_robotti.pdf.

Guerra:2012:BSD

[GR12] Francesco Guerra and Nadia Robotti. La borsa di studio della rivista ‘Missioni’: un punto fermo sulla vicenda di Ettore Majorana. Il Nuovo Saggiatore, 28(1–2):??, February 29, 2012. ISSN 0393-4578 (print), 1827-6148 (elec-

tronic). URL http://static.sif.it:8080/SIF/resources/
public/files/opinioni/op_1202_guerra_robotti-sq.pdf.

Guerra:2013:DDE

[GR13] Francesco Guerra and Nadia Robotti. The disappearance and death of Ettore Majorana. Physics in Perspective (PIP), 15 (2):160–177, June 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://adsabs.harvard.

edu/abs/2013PhP....15..160G;   http://link.springer.
com/article/10.1007/s00016-013-0111-7.

Guerra:2013:PSR

[Gue13] Francesco Guerra. The phenomenon of spontaneous replica symmetry breaking in complex statistical mechanics systems. Journal of Physics: Conference Series, 442(1):012013, 2013. CODEN JPCSDZ. ISSN 1742-6588 (print), 1742-6596 (electronic). URL

http://stacks.iop.org/1742-6596/442/i=1/a=012013.

Guerra:2021:BRE

[Gue21] Francesco Guerra. Book review: Erasmo Recami. The Majorana Case: Letters, Documents, Testimonies, 400 pp., bibl., app. Singapore: World Scientific, 2020. £40 (paper); ISBN 978-981-12-0701-3. Isis, 112(2):422–423, ???? 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hubner:1998:BBS

[HDS^{+}98] K. Hübner, E. Dreisigacker, M. Schreiber, M. G. Huber, N. Straumann, N. Nestle, P. Paufler, B. Eckhard, and H. Schopper. Buch Besprechungen: Sciascia: Das Verschwinden des Ettore Majorana. Calaprice: Einstein sagt. Lindström, Langkau u. Scobel: Physik Kompakt. Bethge: Kernphysik Eine Einführung. Rowan-Robinson: Cosmology. Kimmich: NMR: Tomography, Diffusometry, Relaxometry. Woolfson: An Introduction to X-ray Crystallography. Knauf u Sinai: Classical Nonintegrability, Quantum Chaos. Brown: The Origin of the Concept of Nuclear Forces. (German) [Book reviews: …]. Physikalische Blätter, 54(2):158– 161, 1998. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Heisenberg:1934:CTG

[Hei34] Werner Heisenberg. Considérations théoriques générales sure la structure du noyau. (French) [General theoretical considerations of the structure of the nucleus]. In Cockcroft et al. [CCJ^{+}34], pages 289–335. LCCN ???? Publiés par la commission administrative de l’institut.

Hess:1994:EMS

[Heß94] Georg Heß. Exotic Majorana spinors in (3 + 1)-dimensional space–times. Journal of Mathematical Physics, 35(9):4848–4854, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i9/p4848_s1.

Holstein:2009:MDE

[Hol09] Barry R. Holstein. The mysterious disappearance of Ettore Majorana. Journal of Physics: Conference Series, 173(1):012019, 2009. CODEN JPCSDZ. ISSN 1742-6588 (print), 1742-6596 (electronic). URL http://adsabs.harvard.

edu/abs/2009JPhCS.173a2019H;   http://stacks.iop.org/

1742-6596/173/i=1/a=012019. Carolina International Symposium on Neutrino Physics.

Inguscio:2006:MSF

[Ing06] M. Inguscio. Majorana ‘spin-flip’ and ultra-low temperature atomic physics. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN

???? URL http://pos.sissa.it/archive/conferences/037/
008/EMC2006_008.pdf.

Jackiw:2006:HEMa

[Jac06a] R. Jackiw. Homage to Ettore Majorana. arxiv.org, ??(??):

1–9, October 20, 2006. URL http://adsabs.harvard.edu/
abs/2006hep.th...10228J;   http://arxiv.org/abs/hep-
th/0610228.

Jackiw:2006:HEMb

[Jac06b] R. W. Jackiw. Homage to Ettore Majorana. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN ???? URL http://pos.sissa.it/

archive/conferences/037/009/EMC2006_009.pdf.

Kalogeropoulos:2000:MSU

[Kal00] Nikolaos Kalogeropoulos. Majorana spinors on unoriented surfaces. Journal of Mathematical Physics, 41(5):3113–3124, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kitaev:2006:AES

[Kit06] Alexei Kitaev. Anyons in an exactly solved model and beyond. Annals of Physics, 321(1):2–111, January 2006. CO-DEN APNYA6. ISSN 0003-4916 (print), 1096-035X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0003491605002381.

Klapdor-Kleingrothaus:2002:NMT

[KK02] H. V. Klapdor-Kleingrothaus. Ist das Neutrino ein Majorana-Teilchen?. (German) [Is the neutrino a Majorana particle?]. Physik in unserer Zeit, 33(4):155, 2002. CODEN PHUZAH. ISSN 1521-3943.

Klein:2000:EMG

[Kle00] E. Klein. Ettore Majorana, le génial disparu. (French) [Ettore Majorana, the departed genius]. La Recherche, ??(335):62–63, October 1, 2000. CODEN RCCHBV. ISSN 0029-5671 (print),

1625-9955 (electronic). URL http://www.larecherche.fr/
savoirs/figure-du-passe/ettore-majorana-genial-disparu-
01-10-2000-73252.

Klein:2005:ESF

[Kle05] Etienne Klein. Il était sept fois la révolution: Albert Einstein et les autres . . .. (French) [The seven-times revolution: Albert Einstein and the others …]. Flammarion, Paris, France, 2005. ISBN 2-08-210343-9. 237 pp. LCCN QC16.E5.

Klein:2013:RJP

[Kle13] Etienne Klein. La résurrection d’un jeune physicien. (French) [The resurrection of a young physicist]. Revue de Synthèse, 134 (1):9–27, ???? 2013. ISSN 0035-1776 (print), 1955-2343 (electronic).

Kupsch:1990:EMW

[KT90] J. Kupsch and W. D. Thacker. Euclidean Majorana and Weyl spinors. Fortschritte der Physik = Progress of Physics, 38(1): 35–62, 1990. CODEN FPYKA6. ISSN 1521-3979.

LaLumia:1993:EMA

[LaL93] Joseph LaLumia. Ettore Majorana and the atomic bomb. Professional Press, Chapel Hill, NC, USA, 1993. ISBN 1-880365-67-7. ii + 71 pp. LCCN QC774.M34 L35 1993.

Lanouette:2017:LMM

[Lan17] William Lanouette. The little match’s momentous legacy: The Pope of Physics: Enrico Fermi and the Birth of the Atomic Age by Gino Segrè and Bettina Hoerlin. New York: Henry Holt and Company, 2016, 351 pp. Issues in Science and Technology, 33 (3):93–94, Spring 2017. CODEN ???? ISSN 0748-5492 (print),

1938-1557 (electronic). URL http://issues.org/33-3/book-
review-the-little-matchs-momentous-legacy/.

Lasserre:2007:NOD

[Las07] Thierry Lasserre. Neutrino oscillations — the Double Chooz experiment. Europhysics News, 38(4):20–24, July/August 2007. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL http://www.europhysicsnews.org/articles/
epn/abs/2007/04/epn07401/epn07401.html.

Ledd:2001:TOI

[Led01] Lazarus Ledd. Tra le ombre. (Italian) [Between the shadows]. Edizioni StarComics, 97, 2001.

Licata:2006:ENM

[Lic06] Ignazio Licata. Editorial note: Majorana impact on contemporary physics. Electronic Journal of Theoretical Physics, 3(10):i, ???? 2006. CODEN EJTPBC. ISSN 1729-5254. URL http://

www.ejtp.com/articles/ejtpv3i10p0.pdf. Special Majorana

Centenary Issue: Majorana Legacy in Contemporary Physics.

Levy-Leblond:2004:SPA

[LL04] Jean-Marc Lévy-Leblond. La science et puis après?. (French) [Science, and what next?]. La Recherche, ??(379):111–??, October 1, 2004. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-

9955 (electronic). URL http://www.larecherche.fr/idees/
chroniques/science-puis-apres-01-10-2004-73652.

Majorana:2006:TIT

[MA06] Ettore Majorana and Ennio Arimondo. Theory of the incomplete P′ triplets. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF).

LCCN QC776 .M228 2006. URL http://link.springer.com/
content/pdf/10.1007/978-3-540-48095-2_6.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-540-48095-2/

6/1.pdf. Scientific paper number 5.

Magie:1935:SBP

[Mag35] William Francis Magie. A Source Book in Physics. Source books in the history of the sciences / G. D. Walcott, general ed. Mc-Graw-Hill, New York, NY, USA, 1935. xiv + 620 pp. LCCN ????

Magueijo:2009:BDE

[Mag09] João Magueijo. A brilliant darkness: the extraordinary life and disappearance of Ettore Majorana, the troubled genius of the nuclear age. Basic Books, New York, NY, USA, 2009. ISBN 0-465-00903-4. xxi + 280 pp. LCCN QC774.M34 M35 2009.

Magueijo:2011:GIV

[Mag11] João Magueijo. O grande inquisidor: a vida extraordinária e o desaparecimento misterioso de Ettore Majorana, génio atormentado da era nuclear. (Portuguese) [The Grand Inquisitor: the extraordinary life and mysterious disappearance of Ettore Majorana, tormented genius of the nuclear age], volume 189 of Ciência aberta. Gradiva, Lisboa, Portugal, second edition, 2011. ISBN 989-616-436-3. 454 pp. LCCN ???? 20.50 EUR. URL http:/

/www.gbv.de/dms/casalini/60/60729693.pdf. Translation of

[Mag09] to Portuguese by Helena Ramos.

Magueijo:2016:BRP

[Mag16] João Magueijo. Book review: The Physics of Ettore Majorana: Theoretical, Mathematical, and Phenomenological, Salvatore Esposito, Cambridge U. Press, 2015. $85.00 (382 pp.). ISBN 978-1-107-04402-9. Physics Today, 69(6):61–62, June 2016. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Maiani:2006:MNY

[Mai06] L. Maiani. The Majorana neutrino, 70 years after. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN ???? URL http://pos.

sissa.it/cgi-bin/reader/conf.cgi?confid=37;   http://
www.ct.infn.it/Majorana2006/.

Majorana:1928:TQN

[Maj28] Ettore Majorana. La teoria quantistica dei nuclei radioattivi. (Italian) [The quantum theory of radioactive nuclei]. Master’s thesis., Università di Roma “La Sapienza”, Roma, Italia, July 6, 1928.

Majorana:1929:RUG

[Maj29a] Ettore Majorana. Ricerca di un’espressione generale delle correzioni di Rydberg, valevole per atomi neutri o ionizzati positivamente. (Italian) [Search for a general expression of Rydberg corrections, valid for neutral atoms or positive ions]. Il Nuovo Cimento (8), 6(??):xiv–xvi, ???? 1929. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). Presented on December 29, 1928, during the XXII General Meeting of the Italian Physical Society at the Institute of Physics of the Royal University of Rome, December 28–30, 1928. Also cited in the Jahrbuch über die Fortschritte der Mathematik 55 (1929), 1183.

Majorana:1929:SMN

[Maj29b] Ettore Majorana. Sulla meccanica dei nuclei radioattivi. (Italian) [On the mechanics of radioactive nuclei]. Doctoral thesis, Università di Roma “La Sapienza”, Roma, Italy, July 7, 1929.

Majorana:1931:PTA

[Maj31a] Ettore Majorana. I presunti termini anomali dell’elio. (Italian) [The presumed anomalous terms of helium]. Il Nuovo Cimento (8), 8(??):78–83, ???? 1931. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Majorana:1931:RPF

[Maj31b] Ettore Majorana. Reazione pseudopolare fra atomi di idrogeno. (Italian) [Pseudopolar reaction between hydrogen atoms]. Atti Rendiconti Accademia Nazionale Lincei. Rendiconti, 13(??):58– 61, ???? 1931. Presented by Orso Mario Corbino at the meeting of the Academy on January 4, 1931.

Majorana:1931:SFD

[Maj31c] Ettore Majorana. Sulla formazione dello ione molecolare di elio. (Italian) [On the formation of the helium molecular ion]. Il Nuovo Cimento (8), 8(??):22–28, ???? 1931. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). Announced in Atti Rendiconti Accademia Nazionale Lincei. Rendiconti 12 (1930),

583. Presented by Orso Mario Corbino at the meeting of the Academy on December 7, 1930.

Majorana:1931:TTI

[Maj31d] Ettore Majorana. Teoria dei tripletti P′ incompleti. (Italian) [Theory of incomplete P′ triplets]. Il Nuovo Cimento (8), 8 (1):107–113, March 1931. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Majorana:1932:AOCa

[Maj32a] Ettore Majorana. Atomi orientati in campo magnetico variabile. (Italian) [Oriented atoms in a variable magnetic field]. La Ricerca Scientifica, 3(??):329, ???? 1932. CODEN RISCAZ. ISSN 0035-5011.

Majorana:1932:AOCb

[Maj32b] Ettore Majorana. Atomi orientati in campo magnetico variabile. (Italian) [Oriented atoms in a variable magnetic field]. Il Nuovo Cimento (8), 9(2):43–50, February 1932. CO-DEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (elec-

tronic). URL http://www.nobelprize.org/nobel_prizes/
physics/laureates/1997/;   http://www.nobelprize.org/
nobel_prizes/physics/laureates/2001/;   http://www.nobelprize.
org/nobel_prizes/physics/laureates/2005/.

Majorana:1932:TRP

[Maj32c] Ettore Majorana. Teoria relativistica di particelle con momento intrinseco arbitrario. (Italian) [Relativistic theory of particles with arbitrary intrinsic momentum]. Il Nuovo Cimento (8), 9(10): 335–344, December 1932. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). See comments [Fra66]. That paper reports that both the neutron [Cha32b, Cha32a] and positron [And32] had already been discovered when Majorana’s article was published, but that news had not yet reached Majorana when he was preparing this paper.

Majorana:1933:NRS

[Maj33a] Ettore Majorana. Nuove ricerche sulla teoria dei nuclei. (Italian) [New researches on the theory of nuclei]. La Ricerca Scientifica, 4(1):522–??, ???? 1933. CODEN RISCAZ. ISSN 0035-5011.

Majorana:1933:KGN

[Maj33b] Ettore Majorana. Über die Kerntheorie. (German) [On nuclear theory]. Zeitschrift für Physik, 82(3–4):137–145, March 1933. CODEN ZEPYAA. ISSN 0044-3328. URL http://

adsabs.harvard.edu/abs/1933ZPhy...82..137M;   http://
link.springer.com/article/10.1007/BF01341484. Reprinted

in [Maj65].

Majorana:1937:TSD

[Maj37] Ettore Majorana. Teoria simmetrica dell elettrone e del positrone. (Italian) [Symmetrical theory of the electron and the positron]. Il Nuovo Cimento (8), 14(4):171–184, April 1937. CODEN NU-CIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://en.wikipedia.org/wiki/Majorana_mass;   http://
nl.wikipedia.org/wiki/Majorana-deeltje. In this paper,

Majorana predicted the existence of a new type of particle, now called a Majorana fermion, which is its own antiparticle, and whose existence may have finally been confirmed by experiment seventy years later [FK08, MZF^{+}12, Ser12]. See also comments in [ACM10b]. Esposito [Esp08] reports about this paper “With amazing farsightedness Majorana suggested that the neutrino, which had just been postulated by Wolfgang Pauli and Fermi to explain puzzling features of radioactive beta decay, could be such a particle. This would make the neutrino unique among the elementary particles and, moreover, enable it to have mass. Today many experiments are still devoted to detect these peculiar properties, which include the phenomenon of neutrino oscillations: we have not yet succeeded to find a definite answer to Majorana’s proposal.”.

Majorana:1942:VDL

[Maj42] Ettore Majorana. Il valore delle leggi statistiche nella fisica e nelle scienze sociali. (Italian) [The value of statistical laws in physics and social sciences]. Scientia, 36(??):55–66, ???? 1942. Published posthumously from a manuscript found by Ettore’s brother. Reprinted in English translation in [Man05]. See [Man06b, Man06c] for commentary on the significance of this article.

Majorana:1965:NT

[Maj65] Ettore Majorana. On nuclear theory. In Nuclear Forces [Bri65a], pages 161–169. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965. Reprint of [Maj33b].

Majorana:1987:LAN

[Maj87] Ettore Majorana. Lezioni all’Università di Napoli. (Italian) [Lectures at the University of Naples]. Bibliopolis, Napoli, Italia, 1987. ISBN 88-7088-171-7. 195 pp. Lezioni di fisica teorica che Ettore Majorana tenne all’Università di Napoli nel 1938. Facsimile reproductions of Majorana’s lectures and two letters.

Majorana:2006:SPR

[Maj06] Ettore Majorana. Scattering of an α particle by a radioactive nucleus. Electronic Journal of Theoretical Physics, 3(10):293–303, ???? 2006. CODEN EJTPBC. ISSN 1729-5254. URL http:

//adsabs.harvard.edu/abs/2006EJTP....3j.293M;   http:
//www.ejtp.com/articles/ejtpv3i10p293.pdf. Special Ma-

jorana Centenary Issue: Majorana Legacy in Contemporary Physics. English translation of a previously unpublished paper from 1928; also published in [EMvdMR03].

Maltese:2010:PIE

[Mal10] Giulio Maltese. Il Papa e l’Inquisitore: Enrico Fermi, Ettore Majorana, via Panisperna. (Italian) [The Pope and the Grand Inquisitor: Enrico Fermi, Ettore Majorana, via Panisperna]. Saggi. Zanichelli, Bologna, Italia, 2010. ISBN 88-08-16814-X. 398 pp. LCCN QC774.F4. 27.00 EUR. URL http://www.gbv.de/dms/

casalini/11/11009659.pdf.

Mantegna:2005:PET

[Man05] Rosario Nunzio Mantegna. Presentation of the English translation of Ettore Majorana’s paper: The value of statistical laws in physics and social sciences. Quantitative Finance, 5(2):133– 134, ???? 2005. ISSN 1469-7688 (print), 1469-7696 (electronic). Reprint of [Maj42].

Mantegna:2006:MAB

[Man06a] R. Mantegna. E. Majorana’s article on ‘The value of statistical laws in physics and statistical sciences’. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN ???? URL http://pos.sissa.it/

archive/conferences/037/011/EMC2006_011.pdf.

Mantegna:2006:TAEa

[Man06b] R. Nunzio Mantegna. The tenth article of Ettore Majorana. arxiv.org, ??(??):1–3, August 2006. URL http://

adsabs.harvard.edu/abs/2006physics...8279N;   http://
arxiv.org/abs/physics/0608279. Published in [Man06b].

Mantegna:2006:TAEb

[Man06c] Rosario Nunzio Mantegna. The tenth article of Ettore Majorana. Europhysics News, 37(4):15–17, July/August 2006. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2006ENews.
.37...15M;   http://www.europhysicsnews.org/articles/
epn/abs/2006/04/epn06402/epn06402.html.

March:1975:SCF

[Mar75] Norman H. (Norman Henry) March. Self-Consistent Fields in Atoms: Hartree and Thomas–Fermi Atoms. Pergamon international library of science, technology, engineering and social studies; Selected readings in physics. Pergamon Press, New York, NY, USA; Oxford, UK, 1975. ISBN 0-08-017819-7, 0-08-017820-0 (fleXicover). ix + 233 pp. LCCN QC173 .M3633 1975. See pages 205–213 for an English translation of [Fer28].

Maurizi:2009:DMM

[Mau09] Alessandro Maurizi. Il destino di Majorana: marzo 1938, il genio della fisica atomica Ettore Majorana sparisce nel nulla; non sarà mai ritrovato. (Italian) [The fate of Majorana: March 1938, the genius of atomic physics Ettore Majorana disappears into thin air, never to be found again]. Simple, Macerata, Italia, 2009. ISBN 88-6259-086-5. 199 pp. LCCN QC774.M34. 15.00 EUR.

Majorana:2006:RTP

[MC06] Ettore Majorana and Nicola Cabibbo. Relativistic theory of particles with arbitrary intrinsic angular momentum. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF). LCCN QC776 .M228

2006. URL http://link.springer.com/content/pdf/10.
1007/978-3-540-48095-2_8.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-48095-2/8/1.pdf. Scientific

paper number 7.

Majorana:2006:LFT

[ME06] Ettore Majorana and Salvatore Esposito, editors. Lezioni di fisica teorica. (Italian) [Lectures on theoretical physics], volume 9 of Napoli series on physics and astrophysics. Bibliopolis, Napoli, Italia, 2006. ISBN 88-7088-511-9. xlii + 154 pp.

Meo:2013:LRE

[Meo13] Fioravante Meo. L’ultimo rifugio di Ettore Majorana. (Italian) [The last refuge of Ettore Majorana]. T. Pironti, Napoli, Italia, 2013. ISBN 88-7937-443-5. 101 pp. LCCN QC774.M34 M46 2013. 9.00 EUR.

Majorana:2006:OAV

[MI06] Ettore Majorana and Massimo Inguscio. Oriented atoms in a variable magnetic field. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF).

LCCN QC776 .M228 2006. URL http://link.springer.com/
content/pdf/10.1007/978-3-540-48095-2_7.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-540-48095-2/

7/1.pdf. Scientific paper number 6.

Mignani::1974:ADL

[Mig74] Roberto Mignani. About a Dirac-like equation for the photon according to Ettore Majorana. Lettere al Nuovo Cimento, 11(12): 568–572, November 23, 1974. CODEN NCLTAX. ISSN 0024-1318, 0375-930X, 1827-613X. URL http://link.springer.

com/article/10.1007/BF02812391.

Migneco:2006:HEN

[Mig06] E. Migneco. High energy neutrino astronomy. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN ???? URL http://pos.sissa.it/

archive/conferences/037/012/EMC2006_012.pdf.

Majorana:2006:PAT

[MIM06] Ettore Majorana, Massimo Inguscio, and Francesco Minardi. On the possible anomalous terms of helium. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF). LCCN QC776 .M228

2006. URL http://link.springer.com/content/pdf/10.
1007/978-3-540-48095-2_4.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-48095-2/4/1.pdf. Scientific

paper number 3.

Majorana:2006:STE

[MM06a] Ettore Majorana and Luciano Maiani. A symmetric theory of electrons and positrons. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF).

LCCN QC776 .M228 2006. URL http://link.springer.com/
content/pdf/10.1007/978-3-540-48095-2_10.pdf;   http:
//link.springer.com/content/pdf/bfm:978-3-540-48095-

2/10/1.pdf. Scientific paper number 9.

Majorana:2006:VSL

[MM06b] Ettore Majorana and R. N. Mantegna. The value of statistical laws in physics and social sciences. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF). LCCN QC776

.M228 2006. URL http://link.springer.com/content/
pdf/10.1007/978-3-540-48095-2_11.pdf;   http://link.
springer.com/content/pdf/bfm:978-3-540-48095-2/11/1.

pdf. Scientific paper number 10.

Majorana:2006:PNI

[MPR06] Ettore Majorana, Bruno Preziosi, and Erasmo Recami. Preliminary notes for the inaugural lecture. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF). LCCN QC776 .M228 2006.

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48095-2_12.pdf;   http://link.springer.com/content/
pdf/bfm:978-3-540-48095-2/12/1.pdf. The Inaugural Lecture

(January 1938) at the University of Naples.

Majorana:2006:NT

[MR06] Ettore Majorana and L. A. Radicati di Brozolo. On nuclear theory. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF). LCCN QC776 .M228

2006. URL http://link.springer.com/content/pdf/10.
1007/978-3-540-48095-2_9.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-48095-2/9/1.pdf. Scientific

paper number 8.

Mignani:1975:ADL

[MRB75] R. Mignani, E. Recami, and M. Baldo. About a Dirac-like equation for the photon, according to Ettore Majorana. Technical Report Print-75-0054 (Catania), INFN-ROME-578, ????, Catania, Sicily, Italy, 1975.

Majorana:2006:FMH

[MS06a] Ettore Majorana and Antonio Sasso. On the formation of molecular helium ion. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF).

LCCN QC776 .M228 2006. URL http://link.springer.com/
content/pdf/10.1007/978-3-540-48095-2_3.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-540-48095-2/

3/1.pdf. Scientific paper number 2.

Majorana:2006:PRH

[MS06b] Ettore Majorana and Antonio Sasso. Pseudopolar reaction of hydrogen atoms. In Bassani and the Council of the Italian Physical Society [Bt06], pages xliii + 282. ISBN 3-540-48095-1 (e-book), 3-540-48091-9 (hardcover), 88-7438-031-3 (SIF).

LCCN QC776 .M228 2006. URL http://link.springer.com/
content/pdf/10.1007/978-3-540-48095-2_5.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-540-48095-2/

5/1.pdf. Scientific paper number 4.

Mourik:2012:SMF

[MZF^{+}12] V. Mourik, K. Zuo, S. M. Frolov, S. R. Plissard, E. P. A. M. Bakkers, and L. P. Kouwenhoven. Signatures of Majorana fermions in hybrid superconductor–semiconductor nanowire devices. Science, 336(6084):1003–1007, May 25, 2012. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://arxiv.org/abs/1204.2792;   http://
www.sciencemag.org/content/336/6084/1003.

Naso:2007:EMT

[Nas07] Manuela Naso, editor. Ettore Majorana tra scienza e letteratura: il caso Sciascia; atti del convegno internazionale di studi, Berlino, 21 gennaio 2006. (Italian) [Ettore Majorana between science and literature: the Sciascia case. Proceedings of the international conference, Berlin, January 21, 2006]. Lang, Frankfurt am Main, Germany, 2007. ISBN 3-631-55527-X. LCCN QC774.M34.

Navarro:2011:BRJ

[Nav11] Jaume Navarro. Book review: João Magueijo: A Brilliant Darkness: The Extraordinary Life and Mysterious Disappearance of Ettore Majorana, the Troubled Genius of the Nuclear Age. Isis, 102(2):373–374, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/661691.

Pucci:2006:MAM

[PA06] R. Pucci and G. G. N. Angilella. Majorana: From atomic and molecular, to nuclear physics. Foundations of Physics, 36(10):1554–1572, October 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

adsabs.harvard.edu/abs/2006FoPh...36.1554P;   http://
link.springer.com/article/10.1007/s10701-006-9067-7.

Pavkovic:1970:TMR

[Pav70] Marko I. Pavković. On the two Majorana representations. Journal of Mathematical Physics, 11(4):1123–1125, April 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i4/p1123_s1.

Pessa:2006:MO

[Pes06] Eliano Pessa. The Majorana oscillator. Electronic Journal of Theoretical Physics, 3(10):285–292, ???? 2006. CO-DEN EJTPBC. ISSN 1729-5254. URL http://www.ejtp.com/ articles/ejtpv3i10p285.pdf. Special Majorana Centenary Issue: Majorana Legacy in Contemporary Physics.

Petronzio:2006:NI

[Pet06] R. Petronzio. Neutrinos by INFN. In Rapisarda et al.

[RCC^{+}06a], page ??  LCCN ????  URL http://pos.

sissa.it/cgi-bin/reader/conf.cgi?confid=37;   http://
www.ct.infn.it/Majorana2006/.

PonzdeLeon:1987:STM

[Pon87] Salvo Ponz de Leon. Sulle tracce di Majorana. (Italian) [On the track of Majorana]. TV documentary by Canale Cinque., 1987.

Preziosi:2007:EMI

[PR07] Bruno Preziosi and Erasmo Recami. Ettore Majorana’s inaugural lecture on theoretical physics (Naples, Italy; Jan. 13, 1938). arxiv.org, ??(??):1–34, September 6, 2007. URL http://

adsabs.harvard.edu/abs/2007arXiv0709.0941P;   http://
arxiv.org/abs/0709.0941.

Preziosi:2006:MN

[Pre06] Bruno Preziosi. E. Majorana in Naples. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN ???? URL http://pos.sissa.it/

archive/conferences/037/014/EMC2006_014.pdf.

Pucci:2006:EML

[Puc06] R. Pucci. Ettore Majorana’s legacy to contemporary physics. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN ???? URL http://

pos.sissa.it/archive/conferences/037/015/EMC2006_015.
pdf.

Pauli:1934:QSR

[PW34] W. Pauli and V. Weisskopf. Über die Quantisierung der skalaren relativistischen Wellengleichung. (German) [On the quantization of the scalar relativistic wave equation]. Helvetica Physica Acta, 7(7):709–731, ???? 1934. CODEN HPACAK. ISSN 0018-0238.

Recami:1991:CMC

[RA91] Erasmo Recami and Edoardo Amaldi. Il caso Majorana. Con l’epistolario, documenti, e testimonianze. (Italian) [The Majorana case. With a collection of letters, documents, and evidence].

Mondadori, Milano, Italia, revised and amplified edition, 1991. ISBN 88-04-35344-9. 230 + 16 pp. LCCN QC774.M34.

Recami:2000:CME

[RA00] Erasmo Recami and Edoardo Amaldi. Il caso Majorana: epistolario, documenti, testimonianze. (Italian) [The Majorana case. Collection of letters, documents, evidence], volume 15 of Collana Arcobaleno. Di Renzo, Roma, Italia, second edition, 2000. ISBN 88-8323-028-0. 273 pp. LCCN QC774.M34.

Recami:2002:CMC

[RA02] Erasmo Recami and Edoardo Amaldi. Il caso Majorana. Con l’epistolario, documenti, e testimonianze. (Italian) [The Majorana case. With a collection of letters, documents, and evidence]. Mondadori, Milano, Italia, second edition, 2002. 273 pp. LCCN QC774.M34.

Rapisarda:2006:ICE

[RCC^{+}06a] A. Rapisarda, P. Castorina, F. Catara, E. Migneco, S. Lo Nigro, F. Porto, and E. Rimini, editors. International Conference on Ettore Majorana’s Legacy and the Physics the 21st Century in Commemoration of the Centennial of Majorana’s Birth: EMC2006, October 5–6, 2006, Dipartimento di Fisica e Astronomia, Università di Catania, Italy. ????, ????, 2006. LCCN

???? URL http://pos.sissa.it/cgi-bin/reader/conf.cgi?
confid=37;   http://www.ct.infn.it/Majorana2006/.

Rapisarda:2006:P

[RCC^{+}06b] A. Rapisarda, P. Castorina, F. Catara, E. Migneco, S. Lo Nigro, F. Porto, and E. Rimini. Preface. In International Conference on Ettore Majorana’s Legacy and the Physics the 21st Century in Commemoration of the Centennial of Majorana’s Birth: EMC2006, October 5–6, 2006, Dipartimento di Fisica e Astronomia, Università di Catania, Italy [RCC^{+}06a], page ?? LCCN

???? URL http://pos.sissa.it/archive/conferences/037/
018/EMC2006_018.pdf.

Recami:1998:PMP

[RE98] Erasmo Recami and Myron Evans. Physical meaning of the photon’s wavefunction according to Ettore Majorana. In The enigmatic photon, Vol. 4, volume 90 of Fund. Theories Phys., pages 137–145. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998.

Recami:2007:SML

[RE07]  Erasmo Recami and Salvatore Esposito.  The scientific

manuscripts left unpublished by Ettore Majorana (with outlines of his life and work). arxiv.org, ??(??):1–28, September 8, 2007.

URL http://arxiv.org/abs/0709.1183.

Recami:1975:NDS

[Rec75a] Erasmo Recami. I nuovi documenti sulla scomparsa del fisico Ettore Majorana. (Italian) [New documents on the disappearance of the physicist Ettore Majorana]. Scientia, 110(??):577–583, ???? 1975.

Recami:1975:NED

[Rec75b] Erasmo Recami. New evidence on the disappearance of the physicist Ettore Majorana. Scientia, 110(??):589–598, ???? 1975.

Recami:1978:NEL

[Rec78a] Erasmo Recami. New evidence on the life of Ettore Majorana. In Tachyons, monopoles, and related topics: proceedings of the first session of the interdisciplinary seminars on tachyons and related topics, Erice, 1–15 September, 1976 [Rec78b], pages 269–277 + 1. ISBN 0-444-85165-8. LCCN QC793.5.T322 T3.

Recami:1978:TMR

[Rec78b] Erasmo Recami, editor. Tachyons, monopoles, and related topics: proceedings of the first session of the interdisciplinary seminars on tachyons and related topics, Erice, 1–15 September, 1976. North-Holland, Amsterdam, The Netherlands, 1978. ISBN 0-444-85165-8. LCCN QC793.5.T322 T3.

Recami:1987:CMC

[Rec87] Erasmo Recami. Il caso Majorana. Con l’epistolario, documenti, e testimonianze. (Italian) [The Majorana case. With a collection of letters, documents, and evidence]. Mondadori, Milano, Italia, 1987. ISBN 88-04-29895-2. xii + 175 + 16 pp. LCCN QC774.M34.

Recami:1998:CSM

[Rec98] Erasmo Recami. Catalog of the scientific manuscripts left by Ettore Majorana (with a recollection of E. Majorana, sixty years after his disappearance). arxiv.org, ??(??):1–81, December

1998. URL http://adsabs.harvard.edu/abs/1998physics..
10023R;   http://arxiv.org/abs/physics/9810023. A pre-

liminary version of this article appears in [Rec99a].

Recami:1999:CSM

[Rec99a] Erasmo Recami. Catalog of the scientific manuscripts left by Ettore Majorana (with a recollection of E. Majorana sixty years after his disappearance). Quaderni di Storia della fisica, 5(??): 19–68, ???? 1999. ISSN 1594-9974 (print), 1827-6164 (electronic).

Recami:1999:EMO

[Rec99b] Erasmo Recami. Ettore Majorana: l’opera scientifica edita e inedita. (Italian) [Ettore Majorana: published and unpublished scientific work]. Il Nuovo Saggiatore, 15(3–4):14–29, May/August 1999. CODEN ???? ISSN 0393-4578 (print), 1827-6148 (electronic).

Recami:1999:REM

[Rec99c] Erasmo Recami. Ricordo di Ettore Majorana a sessant’anni dalla sua scomparsa: L’opera scientifica edita ed inedita. (Italian) [I remember Ettore Majorana sixty years after his death: published and unpublished scientific work]. Quaderni di Storia della fisica, 5(??):??, ???? 1999. ISSN 1594-9974 (print), 1827-6164 (electronic).

Recami:2006:MHH

[Rec06a] E. Recami. E. Majorana: his human and scientific personality. In Rapisarda et al. [RCC^{+}06a], page ?? LCCN

???? URL http://pos.sissa.it/archive/conferences/037/
016/EMC2006_016.pdf.

Recami:2006:SWE

[Rec06b] Erasmo Recami. The scientific work of Ettore Majorana: an introduction. Electronic Journal of Theoretical Physics, 3(10):1–10, ???? 2006. CODEN EJTPBC. ISSN 1729-5254. URL http://

adsabs.harvard.edu/abs/2006EJTP....3j...1R;   http://
www.ejtp.com/articles/ejtpv3i10p1.pdf. Special Majorana

Centenary Issue: Majorana Legacy in Contemporary Physics.

Recami:2007:EMS

[Rec07a] Erasmo Recami. Ettore Majorana’s scientific (and human) personality. arxiv.org, ??(??):1–12, August 21, 2007. URL http://

arxiv.org/abs/0708.2855.

Recami:2007:LEM

[Rec07b] Erasmo Recami. L’articolo di Ettore Majorana su ‘Il valore delle Leggi Statistiche nella Fisica e nelle Scienze Sociali’. (Italian)

[Ettore Majorana’s article on ‘The value of Statistical Laws in Physics and in Social Sciences’]. arxiv.org, ??(??):1–21, Septem-

ber 21, 2007. URL http://arxiv.org/abs/0709.3537.

Recami:2008:CME

[Rec08] Erasmo Recami. Il caso Majorana: epistolario, documenti, testimonianze. (Italian) [The Majorana case. Collection of letters, documents, evidence]. Di Renzo, Roma, Italia, third edition, 2008. ISBN 88-8323-028-0. 273 + 8 pp. LCCN QC774.M34.

Recami:2009:EMI

[Rec09] Erasmo Recami. Ettore Majorana: Introduzione a vita e opere. (Italian) [Ettore Majorana: An introduction to his life and work]. arxiv.org, ??(??):1–56, November 4, 2009. URL http:

//adsabs.harvard.edu/abs/2009arXiv0911.0933R;   http:
//arxiv.org/abs/0911.0933.

Recami:2010:ML

[Rec10] Erasmo Recami. Majorana’s legacy. Physics World, 23(4): 20, April 2010. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/23/04/phwv23i04a28.pdf.

Recami:2011:CME

[Rec11] Erasmo Recami. Il caso Majorana: epistolario, documenti, testimonianze. (Italian) [The Majorana case. Collection of letters, documents, evidence]. Di Renzo, Roma, Italia, fourth edition, 2011. ISBN 88-8323-265-8. 273 + 8 pp. LCCN QC774.M34.

Recami:2020:MCL

[Rec20] Erasmo Recami. The Majorana Case: letters, documents, testimonies. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2020. ISBN 981-12-0701-1 (hardcover), 981-12-0769-0 (paperback). xvi + 356 pp. LCCN QC774.M34 A413 2020.

Reich:2012:SCM

[Rei12] Eugenie Samuel Reich. A solid case for Majorana fermions: First compelling evidence of self-annihilating entities in a semiconductor is a step forward for quantum computing. Nature, 483(7388): 132, March 8, 2012. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/news/a-
solid-case-for-majorana-fermions-1.10174.

Rekami:1983:SND

[Rek83] Èrazmo Rekami. Some new data on the life and activity of Ettore Majorana. In Problems in statistical and quantum physics, pages 3–22. Univ. Druzhby Narodov, Moscow, USSR, 1983.

Rathsman:1997:MMC

[RI97] J. Rathsman and G. Ingelman. MAJOR 1.5 — a Monte Carlo generator for heavy Majorana neutrinos in e–p collisions. Computer Physics Communications, 101(1–2):143–154, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001324.

Rokhinson:2012:FCJ

[RLF12] Leonid P. Rokhinson, Xinyu Liu, and Jacek K. Furdyna. The fractional a.c. Josephson effect in a semiconductor superconductor nanowire as a signature of Majorana particles. Nature Physics, 8(11):795–799, September 23, 2012. CO-DEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (elec-

tronic). URL http://www.nature.com/nphys/journal/v8/
n11/full/nphys2429.html.

Rogers:2013:NDY

[Rog13] J. D. Rogers. The neutron’s discovery — 80 years on. Physics Procedia, 43:1–9, 2013. CODEN PPHRCK. ISSN 1875-

3892. URL http://adsabs.harvard.edu/abs/2013PhPro..43.
...1R.

Roncoroni:2011:PTN

[Ron11] Stefano Roncoroni. il promemoria ‘Tunisi’: un nuovo tassello del caso Majorana. (Italian) [The reminder of ‘Tunis’: a new piece in the Majorana case]. Il Nuovo Saggiatore, 27(5–6): 58–69, ???? 2011. ISSN 0393-4578 (print), 1827-6148 (elec-

tronic). URL http://prometeo.sif.it/papers/online/sag/
027/05-06/pdf/08-ilnostromondo.pdf.

Roncoroni:2013:EMS

[Ron13] Stefano Roncoroni. Ettore Majorana, lo scomparso e la decisione irrevocabile. (Italian) [Ettore Majorana, the deceased and the irrevocable decision]. Navigazioni. Editori internazionali riuniti, Roma, Italia, 2013. ISBN 88-359-9237-0. 415 pp. LCCN

QC774.M34 R66 2013. URL http://www.sif.it/attivita/
saggiatore/recensioni/roncoroni.

Recami:2007:ESM

[RRR07] Erasmo Umberto M. Recami, Umberto Victor G. Recami, and Erasmo Recami. Einstein, Sciascia, Majorana, Amaldi, e il rapporto tra intellettuali e potere. (Italian) [Einstein, Sciascia, Majorana, Amaldi, and the relationship of intellectuals with the power]. arxiv.org, ??(??):1–53, September 28, 2007. URL http:

//arxiv.org/abs/0709.4672.

Rigden:2009:BNA

[RS09] John S. Rigden and Roger H. Stuewer. Book notes [Alan Hirshfeld, Eureka Man: The Life and Legacy of Archimedes (New York: Walker and Company, 2009, viii + 228 pages); Francesco Guerra and Nadia Robotti, Ettore Majorana: Aspects of His Scientific and Academic Activity (Pisa: Edizioni Della Normale, 2008, xii + 243 pages)]. Physics in Perspective (PIP), 11(4): 454–456, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0435-5.

Russo:1990:EMG

[Rus90] Bruno Russo. Ettore Majorana: un giorno di marzo. (Italian) [Ettore Majorana: a day in March]. TV documentary produced by Rai III Sicilia., 1990.

Russo:1997:EMG

[Rus97] Bruno Russo. Ettore Majorana: un giorno di marzo. (Italian) [Ettore Majorana: a day in March], volume 19 of Saggi. Flaccovio, Palermo, Italia, 1997. ISBN 88-7804-137-8. 131 pp. LCCN QC16.M335 R87 1997.

Schmauss:1903:HMG

[Sch03] August Schmauss. Über die von Hrn. Majorana gefundene Doppelbrechung im magnetischen Felde. (German) [On Mr. Majorana’s detection of birefringence in a magnetic field]. Annalen der Physik (1900), 317(9):186–195, 1903. ISSN 1521-3889.

Schwinger:1977:MF

[Sch77] Julian Schwinger. The Majorana formula. Transactions of the New York Academy of Sciences, Series II, 38(1 Series II):170–184, 1977. CODEN TNYAAE. ISSN 2164-0947.

Sciascia:1975:SMI

[Sci75] Leonardo Sciascia. La scomparsa di Majorana. (Italian) [The disappearance of Majorana], volume 139 of Nuovi coralli. Einaudi, Torino, Italai, 1975. ISBN 88-06-43737-2. iv + 79 pp.

Sciascia:1977:PSM

[Sci77a] Leonardo Sciascia. I pugnalatori. La scomparsa di Majorana. (Italian) [The slashers. The disappearance of Majorana]. Club degli editori, Milano, Italia, 1977. ISBN ???? 178 pp. LCCN ????

Sciascia:1977:MFS

[Sci77b] Leonardo Sciascia. Majoranas försvinnande. (Swedish) [The disappearance of Majorana]. PAN/Norstedt, Stockholm, Sweden, 1977. ISBN 91-1-771581-4. 78 pp. LCCN ???? Translation to Swedish by Sören Johansson of [Sci75].

Sciascia:1978:NDM

[Sci78] Leonardo Sciascia. Los navajeros. La desaparición de Majorana. (Spanish) [The slashers. The disappearance of Majorana]. Noguer, Barcelona, España, 1978. ISBN 84-279-1172-6. 221 pp. LCCN ???? Spanish translation by Javier Villalba of I pugnalatori / La scomparsa di Majorana [Sci77a].

Sciascia:1980:FMG

[Sci80] Leonardo Sciascia. Der Fall Majorana. (German) [The case of Majorana]. Frankfurt am Main, West Germany, Ullstein, 1980. ISBN 3-548-20062-1. 122 pp. LCCN ???? Translation to German by Ruth Wright and Ingeborg Brandt of [Sci75].

Sciascia:1982:SMI

[Sci82] Leonardo Sciascia. La scomparsa di Majorana. (Italian) [The disappearance of Majorana]. Einaudi, Torino, Italai, third edition, 1982. ISBN ???? ???? pp.

Sciascia:1987:MAD

[Sci87] Leonardo Sciascia. The Moro affair; and, the disappearance of Majorana. Carcanet, Manchester, UK, 1987. ISBN 0-85635-700-6. 220 pp. LCCN DG579.M63 S3613 1987. English translations by Sacha Rabinovitch of L’affaire Moro and La scomparsa di Majorana.

Sciascia:1996:VEM

[Sci96] Leonardo Sciascia. Das Verschwinden des Ettore Majorana. (German) [The Disappearance of Ettore Majorana]. Elster Verlag, Baden-Baden, Germany, 1996. ISBN 3-89151-235-X. 97 pp. LCCN 0.2sci a0160.

Sciascia:1997:SMI

[Sci97] Leonardo Sciascia. La scomparsa di Majorana. (Italian) [The disappearance of Majorana]. Adelphi Edizione, Milano, Italia, 1997. ISBN 88-459-1269-8. 119 pp. LCCN ???? With an essay by Lea Ritter Santini.

Sciascia:2001:SMI

[Sci01] Leonardo Sciascia. La scomparsa di Majorana. (Italian) [The disappearance of Majorana]. Mondolibri, Milano, Italia, 2001. ISBN ???? 119 pp. LCCN ???? With an essay by Lea Ritter Santini.

Sciascia:2006:SMI

[Sci06] Leonardo Sciascia. La scomparsa di Majorana. (Italian) [The disappearance of Majorana], volume 249 of Gli Adelphi. Adelphi, Milano, Italia, sixth edition, 2006. ISBN 88-459-1871-8. 119 pp. LCCN ???? With an essay by Lea Ritter Santini.

Sciascia:2012:DMF

[Sci12] Leonardo Sciascia. La disparition de Majorana. (French) [The disappearance of Majorana]. Éd. Allia, Paris, France, 2012. ISBN 2-84485-431-1. 110 pp. LCCN ???? French translation by Mario Fusco of [Sci75].

Segre:1970:EFPb

[Seg70] Emilio Segrè. Enrico Fermi: physicist. University of Chicago Press, Chicago, IL, USA, 1970. ISBN 0-226-74472-8. x + 276 pp. LCCN QC16.F46 S4.

Segre:1988:LIE

[Seg88a] E. Segrè. Una lettera inedita di Ettore Majorana. . (Italian) [an unpublished letter of Ettore Majorana]. Storia Contemporanea, 19(1):107–111, ???? 1988. ISSN 0039-1875.

Segre:1988:FSR

[Seg88b] Emilio Segrè. The Fermi school in Rome. European Journal of Physics, 9(2):83–87, April 1988. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://iopscience.

iop.org/0143-0807/9/2/001.

Segre:2010:BRJ

[Seg10] Gino Segrè. Book review: João Magueijo, A Brilliant Darkness: The Extraordinary Life and Mysterious Disappearance of Ettore Majorana, the Troubled Genius of the Nuclear Age. New York: Basic Books, 2009. Physics in Perspective (PIP), 12(3):365–367, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Service:2012:PDN

[Ser12] Robert F. Service. Physicists discover new type of particle—sort of. Science Now, ??(??):??, April 12, 2012. URL http://news.

sciencemag.org/sciencenow/2012/04/majorana-fermions-
found.html.

Segre:2016:PPE

[SH16] Gino Segrè and Bettina Hoerlin. The Pope of Physics: Enrico Fermi and the birth of the atomic age. Henry Holt and Company, New York, NY, USA, 2016. ISBN 1-62779-005-5, 1-62779-006-3 (e-book). xi + 351 pp. LCCN QC16.F46 S44 2016.

Shchepkin:1986:MND

[Shc86] M. G. Shchepkin. Majorana neutrino and double beta decay. Physics-Uspekhi, 29(11):1062–1063, November 1986. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL

http://ufn.ru/en/articles/1986/11/g/.

Simpao:2006:ESM

[Sim06] Valentino A. Simpao. Exact solution of Majorana equation via Heaviside operational ansatz. Electronic Journal of Theoretical Physics, 3(10):239–247, ???? 2006. CODEN EJTPBC. ISSN 1729-5254. URL http://www.ejtp.com/ articles/ejtpv3i10p239.pdf. Special Majorana Centenary Issue: Majorana Legacy in Contemporary Physics.

Slaughter:2010:BRJ

[Sla10] Aimee C. E. Slaughter. Book review: João Magueijo, A Brilliant Darkness: The Extraordinary Life and Mysterious Disappearance of Ettore Majorana, the Troubled Genius of the Nuclear Age. American Journal of Physics, 78(4):446–447, April 2010. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2010AmJPh..
78..446M;   http://link.aip.org/link/?AJP/78/446/1.

Sommerfeld:1932:IAD

[Som32] Arnold Sommerfeld. Integrazione asintotica dell’equazione differenziale di Thomas–Fermi. (Italian) [Asymptotic integration of the Thomas–Fermi differential equation]. Rendiconti dell’Accademia Nazionale dei Lincei, 15(??):293–308, ???? 1932. CODEN AANLAW. ISSN 0001-4435.

Spruch:1991:PNT

[Spr91] Larry Spruch. Pedagogic notes on Thomas–Fermi theory (and on some improvements): atoms, stars, and the stability of bulk matter. Reviews of Modern Physics, 63(1):151–209, January 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.63.151;   http://rmp.aps.org/abstract/RMP/
v63/i1/p151_1.

Sau:2012:RRP

[SS12] Jay D. Sau and S. Das Sarma. Realizing a robust practical Majorana chain in a quantum-dot-superconductor linear array. Nature Communications, 3(7):964, July 17, 2012. CODEN NCAOBW. ISSN 2041-1723 (electronic). URL http://www.nature.com/

ncomms/journal/v3/n7/full/ncomms1966.html.

Stoyanov:1968:MRL

[ST68] D. Tz. Stoyanov and I. T. Todorov. Majorana representations of the Lorentz group and infinite-component fields. Journal of Mathematical Physics, 9(12):2146–2167, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i12/p2146_s1.

Stajic:2017:SMM

[Sta17] Jelena Stajic. Sighting of magnetic Majorana fermions? Science, 356(6342):1040, 2017. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://science.

sciencemag.org/content/356/6342/1040.2.

Stepanovsky:2002:EMM

[Ste02] Y. P. Stepanovsky. Ettore Majorana and Matvei Bronstein (1906–1938): Men and scientists. In Bergmann and De Sabbata [BD02], pages 435–458. ISBN 1-4020-0592-X. LCCN QC178 .A38

2002. URL http://adsabs.harvard.edu/abs/2002aibq.conf.
.435S.

Sullivan:1987:FRM

[Sul87] Walter Sullivan. Finding on radioactivity may upset physics law: Rare form of decay believed to be found in mine. New York Times, ??(??):A14, January 14, 1987. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/110629808/13F2FFCB7B870C83A2B/2.

Sutton:1992:SN

[Sut92] Christine Sutton. Spaceship Neutrino. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-36404-3 (hardcover), 0-521-36703-4 (paperback). xiv + 244 pp. LCCN QC793.5.N42 S88 1992. URL http://www.loc.

gov/catdir/description/cam021/92004215.html;   http:
//www.loc.gov/catdir/samples/cam034/92004215.html;
http://www.loc.gov/catdir/toc/cam021/92004215.html.

Sciascia:2011:VEM

[SW11] Leonardo Sciascia and Ruth Wright. Das Verschwinden des Ettore Majorana. (German) [The Disappearance of Ettore Majorana], volume 652 of Wagenbachs Taschenbuch. Wagenbach, Berlin, Germany, 2011. ISBN 3-8031-2652-5. 90 pp. LCCN ????

Thomas:1927:CAF

[Tho27] Llewellyn H. Thomas. The calculation of atomic fields. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 23(5):542–548, January 1927. CODEN PCPSA4. ISSN 0008-1981. This is the Thomas side of the Thomas–Fermi (or Fermi–Thomas) model. See [Fer27] for the Fermi side.

Tolksdorf:2010:RBM

[Tol10] Jürgen Tolksdorf. Real Z_{2}-bigradings, Majorana modules, and the standard model action. Journal of Mathematical Physics, 51(5):053525, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i5/p053525_s1.

Turchetti:2008:BRV

[Tur08] Simone Turchetti. Book review: Valeria Del Gamba, Il Ragazzo di Via Panisperna. l’avventurosa vita del fisico Franco Rasetti. Torino: Bollati Boringhieri, 2007. 167 pp., ISBN 978-88-339-1746-7. Nuncius, 23(1):174–175, ???? 2008. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00427.

Varlamov:2002:AAF

[Var02] V. V. Varlamov. About algebraic foundations of Majorana– Oppenheimer quantum electrodynamics and de Broglie–Jordan neutrino theory of light. Annales de la Fondation Louis de Broglie, 27(2):273–286, 2002. ISSN 0182-4295.

vonBorzeszkowski:1985:MER

[vBT85] H.-H. von Borzeszkowski and H. J. Treder. On the Majorana equations and the rest mass of neutrinos. Foundations of Physics, 15(2):193–199, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00735291.

vanCalmthout:2017:RQSc

[vCK17] Martijn van Calmthout and Leo P. Kouwenhoven. Real Quanta: Simplifying Quantum Physics for Einstein and Bohr. Dundurn Press Limited, Toronto, ON, Canada, 2017. ISBN 1-4597-4049-1 (paperback), 1-4597-4050-5 (PDF), 1-4597-4051-3 (epub). 192 pp. LCCN QC174.123 .C3513 2017.

Vittorini-Orgeas:2009:MTA

[VOB09] Alessandra Vittorini-Orgeas and Antonio Bianconi. From Majorana theory of atomic autoionization to Feshbach resonances in high temperature superconductors. Journal of Superconductivity and Novel Magnetism, 22(3):215–221, April 2009. CO-DEN JSNMBN. ISSN 1557-1939 (print), 1557-1947 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10948-008-0433-x. See [Maj31d].

Weaver:1976:MT

[Wea76] D. L. Weaver. On the Majorana transformation. Journal of Mathematical Physics, 17(4):485–489, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i4/p485_s1.

Weaver:1987:BRM

[Wea87] William Weaver. Book review: Murder in the people’s prison: The Moro Affair and The Mystery of Majorana by Leonardo Sciascia. Translated by Sacha Rabinovitch. 175pp. New York: Carcanet. $16.95. New York Times, ??(??):BR10, April 5, 1987. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/110802650/13F2FFCB7B870C83A2B/1.

Weisskopf:1985:PMP

[Wei85] Victor F. Weisskopf. Personal memories of Pauli. Physics Today, 38(12):36–41, December 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v38/i12/p36/s1.

Williams:2012:MFD

[WGG12] James R. Williams and David Goldhaber-Gordon. Majorana fermions: Doubling down on Majorana. Nature Physics, 8(11): 778–779, November 6, 2012. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic). URL http://www.nature.com/

nphys/journal/v8/n11/full/nphys2459.html.

Wick:1981:X

[Wic81] Gian Carlo Wick. xxx. Atti Accad. Sci. Torino Suppl. 2, 115(??):

13–??, ???? 1981. URL http://www.nap.edu/readingroom.
php?book=biomems&page=gwick.html.

Wigner:1939:URI

[Wig39] E. P. Wigner. On unitary representations of the inhomogeneous Lorentz group. Annals of Mathematics (2), 40(1):149–204, January 1939. CODEN ANMAAH. ISSN 0003-486X (print),

1939-8980 (electronic). URL http://www.jstor.org/stable/

1968551. Reprinted in [Wig89].

Wigner:1989:URI

[Wig89] E. Wigner. On unitary representations of the inhomogeneous Lorentz group. Nuclear Physics B, Proceedings Supplements, 6(1):9–64, March 1989. CODEN NPBSE7. ISSN 0920-5632 (print), 1873-3832 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0920563289904027.

Reprint of [Wig39].

Wilczek:2009:MR

[Wil09] Frank Wilczek. Majorana returns. Nature Physics, 5(9):614–618, September 2009. CODEN NPAHAX. ISSN 1745-2473 (print),

1745-2481 (electronic). URL http://www.nature.com/nphys/
journal/v5/n9/full/nphys1380.html.

Wilczek:2012:QPM

[Wil12] Frank Wilczek. Quantum physics: Majorana modes materialize. Nature, 486(7402):195–197, June 14, 2012. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v486/
n7402/full/486195a.html.

Zaslavskii:2006:EMQa

[Zas06a] O. B. Zaslavskii. Ettore Majorana: Quantum mechanics of destiny. arxiv.org, ??(??):1–16, April 2006. URL http://adsabs.

harvard.edu/abs/2006physics...5001Z. To appear in Priroda.

Zaslavskii:2006:EMQb

[Zas06b] O. B. Zaslavskii. Ettore Majorana: Quantum mechanics of destiny. Priroda, 11(??):55–63, 2006.

Zichichi:1966:SWI

[Zic66] Antonino Zichichi, editor. Strong and weak interactions: present problems. 1966 International School of Physics “Ettore Majorana”, a CERN-MPI-NATO Advanced Study Institute, Erice, June 19–July 4. Academic Press, New York, USA, 1966. LCCN ????

Zichichi:2006:EMGa

[Zic06a] Antonino Zichichi. Ettore Majorana: genius and mystery. CERN Courier, 46(6):23–26, July/August 25, 2006. URL http://

cerncourier.com/cws/article/cern/29664.

Zichichi:2006:EMGb

[Zic06b] Antonino Zichichi. Ettore Majorana: Genius and mystery. Report, Ettore Majorana Foundation and Centre for Scientific Culture (EMFCSC), Erice, Sicily, Italy, September 7, 2006.

62 pp. URL http://www.ccsem.infn.it/em/EM_genius_and_
mystery.pdf.

Zichichi:2006:MGM

[Zic06c] Antonino Zichichi. Ettore Majorana: Genius and mystery. In Rapisarda et al. [RCC^{+}06a], pages 1–47. LCCN ????

URL http://pos.sissa.it/archive/conferences/037/017/
EMC2006_017.pdf.

Zichichi:2007:EMH

[Zic07] Antonino Zichichi. Ettore Majorana: His genius and long-lasting legacy: 1906–2006. Società Italiana di Fisica, Bologna, Italia, 2007. ISBN ???? 143 pp. LCCN ????

Ziebolz:2000:GLE

[Zie00] Burkhard Ziebolz. Das geheime Leben des Ettore Majorana. (German) [The secret life of Ettore Majorana]. KBV-Roman. KBV-Verl.- und Medienges, Hamm, Germany, 2000. ISBN 3-934638-53-8 (paperback). 283 pp. LCCN ????

Xing:2006:WDM

[zXZ06] Zhi zhong Xing and He Zhang. Why do Majorana neutrinos run faster than Dirac neutrinos? Electronic Journal of Theoretical Physics, 3(10):191–209, ???? 2006. CO-DEN EJTPBC. ISSN 1729-5254. URL http://www.ejtp.com/ articles/ejtpv3i10p191.pdf. Special Majorana Centenary Issue: Majorana Legacy in Contemporary Physics.